### v1

# Asymmetry Analysis of Bilateral Shapes

Abstract
Many biological objects possess bilateral symmetry about a midline or midplane, up to a “noise” term. This paper uses landmark-based methods to measure departures from bilateral symmetry, especially for the two-group problem where one group is more asymmetric than the other. In this paper, we formulate our work in the framework of size-and-shape analysis including registration via rigid body motion. Our starting point is a vector of elementary asymmetry features defined at the individual landmark coordinates for each object. We introduce two approaches for testing. In the first, the elementary features are combined into a scalar composite asymmetry measure for each object. Then standard univariate tests can be used to compare the two groups. In the second approach, a univariate test statistic is constructed for each elementary feature. The maximum of these statistics lead to an overall test statistic to compare the two groups and we then provide a technique to extract the important features from the landmark data. Our methodology is illustrated on a pre-registered smile dataset collected to assess the success of cleft lip surgery on human subjects. The asymmetry in a group of cleft lip subjects is compared to a group of normal subjects, and statistically significant differences have been found by univariate tests in the first approach. Further, our feature extraction method leads to an anatomically plausible set of landmarks for medical applications.
Keywords— asymmetry scores, cleft lip disfigurement, facial trauma surgery, feature selection, registration, shape analysis, size-and-shape, smile.
1 Introduction
Bilateral symmetry (left-right symmetry) is a key property in many biological settings. An object (treated as a set of unlabelled points known as landmarks) is said to be bilaterally symmetric if its reflection through some midline or midplane $\mathcal{P}$ is exactly the same as the original object. In practice, exact bilateral symmetry seldom holds and it is of interest to study the extent of any asymmetry. Research about the symmetries of objects in the real world has a long history (see, for example, Weyl, 1952). Of particular interest is the hypothesis of bilateral symmetry, especially in biology (Palmer and Strobeck, 1986). Statistical methods to test asymmetry were first formalized in Mardia et al. (2000). Subsequently, a wide variety of statistical methods have been developed; see, for examples, Kent and Mardia (2001), Bock and Bowman (2006), Ajmera et al. (2022) and Ajmera et al. (2023).
The focus of this paper is on landmark-based objects as in Mardia et al. (2000). Each object is a set of $K$ labelled points or landmarks, represented as a $K\times M$ configuration matrix giving the positions of $K$ landmarks in $M$ dimensions. The cases $M=2$ and $M=3$ are the most important in practice. Call an object bilateral if its landmarks can be divided into two categories: pairs and solos, where paired landmarks locate on both sides of $\mathcal{P}$, while solos are unpaired. For a bilaterally symmetric object the two landmarks in a pair are equally spaced about the midplane and the solos lie on the midplane. If there are $K_{P}$ pairs of landmarks and $K_{S}$ solo landmarks, then the total number of landmarks is $K=2K_{P}+K_{S}$. If the dataset has not been pre-registered, then registration is required. The details of registration are given in Section 2.1.
The asymmetry information in an object can be represented as a vector of coordinatewise elementary asymmetry features, $\bm{d}$, say, defined in more details in Section 2.2. Increasing the magnitude of any element of this vector indicates greater asymmetry. The elementary feature vector is also the starting point of Bock and Bowman (2006) and Patel et al. (2023) but we define its elements explicitly in Section 2.2.
Given two groups of objects, a natural question is whether one group is more asymmetric than the other. Section 3 looks at two testing strategies. In the first strategy the elements of the feature vector $\bm{d}$ are combined into a single number called a composite score. Then a univariate test such as the two-sample t-test or the Mann-Whitney U test can be used to compare the two groups. The use of a univariate summary statistic for each object facilitates the comparison between objects and between groups of objects. The mathematical details are given in Section 3.1, our composite score extends the asymmetry measures used by Bock and Bowman (2006) and Patel et al. (2023).
In the second strategy, the two groups are compared using each of the elements in $\bm{d}$, yielding a collection of test statistics. The maximum of these test statistics gives an overall test statistic that can be used to test for a difference between the two groups. This approach is similar to the union-intersection test (UIT) in multivariate analysis, see, for example, Mardia et al. (2024). In particular, when the null hypothesis is rejected, it is possible to investigate which elementary features are responsible. The mathematical details are given in Section 3.2.
The motivation for the paper is given by a set of smile data on human subjects. Each subject was asked to smile, and a four-dimensional movie was made of the face (three spatial dimensions plus time using Di4D motion capture system). A set of landmarks was identified around the lips and followed through the frames of the movie. The subjects come from two groups: normal subjects and subjects who have had cleft lip surgery. The objective is to assess how successful the surgery has been; that is, are the cleft lip subjects any more asymmetric than normal people. The details of the smile data are provided in Section 4. The data is pre-registered and the objects at three fixed frames (the beginning, middle and end of the smile) are analyzed in Section 4 following Patel et al. (2023). We end the paper with some discussion in Section 5. The procedures for estimating the midline or midplane are given in Appendix A.
2 Describing asymmetry
2.1 Registration
We will introduce the following notations: let $X$, with elements $X[k,m]$, be a $K\times M$ configuration matrix giving the positions of $K$ landmarks in $M$ dimensions for a single object or subject or individual. We write as $X[k,\ ]\in\mathbb{R}^{M}$ for the coordinates of $k$th landmark, $k=1,\dots,K$. The most important choice for $M$ is $M=3$, and this case is emphasized in the presentation, but the mathematical theory is valid for any $M\geq 1$. In general, upper case letters are used for matrices and bold lower case letters for vectors. In addition the letters $K,M,J$ are reserved for the number of landmarks, number of dimensions, and number of elementary features (Section 2.2), respectively, with typical indices given by the lower case letters $k,m,j$. In general, each data object $X$ will be viewed as a noisy version of a bilaterally symmetric configuration.
Rigid body transformations If $X$ is a configuration, then a rigid body motion takes $X$ to $\bm{1}_{M}\bm{c}^{T}+XR=X^{*}$, say, where $\bm{c}$ is an $M$-dimensional translation vector and $R$ is an $M\times M$ rotation matrix. For the purposes of this paper, we say that $X$ and $X^{*}$ have the same size-and-shape following the standard terminology in shape analysis (see, for example, Dryden and Mardia (2016)). That is, the size-and-shape of $X$ is the equivalence class of configurations under the group of rigid body motions. The asymmetry information carried by the size-and-shape is of interest and such information is unaffected under rigid body motion. The size is emphasized here, since the size information is taken into consideration. Recall that the term “shape” is a standard terminology in statistical shape analysis which means the equivalence class under the larger group of similarity transformations so the scale is also filtered out.
A hyperplane $\mathcal{P}$ in $\mathbb{R}^{M}$ can be written in the form $\mathcal{P}=\{\bm{x}\in\mathbb{R}^{M}:\bm{n}^{T}\bm{x}=b\}$, in terms of a unit normal vector $\bm{n}$, say, and a scalar offset term $b$, say. When $M=3$ a hyperplane becomes a two-dimensional plane, and when $M=2$ a hyperplane becomes a one-dimensional line. For simplicity we refer to hyperplanes as planes everywhere. Under a rigid body motion, $\bm{n}$ is transformed to $R^{T}\bm{n}$ and $b$ is transformed to $b+\bm{n}^{T}R\bm{c}$.
Two types of Registration For every bilateral data object $X$, suppose there is an associated midplane $\mathcal{P}$. Ways in which $\mathcal{P}$ might be determined are disussed below. For data analysis, it is helpful to “register” $X$ using a rigid body motion so that after transformation $\bm{n}=\bm{e}_{1}$ and $b=0$, i.e., the midplane passes through the origin and the normal direction is given by the first coordinate axis, while still keeps the shape of $X$ unchanged.
We will distinguish two types of registration:
  1. 1.
    Axis registration
  2. 2.
    Basis registration.
Axis registration takes the form described in the previous paragraph. In terms of the human head, after registration, the left-right axis lies along the first coordinate axis. However, the position of the down-up and back-front axes are not clearly distinguished from one another. It is possible to tilt the head forwards or backwards while still preserving axis registration.
In some settings it is possible to carry out a stronger version of registration, we will call basis registration, in which all the coordinate axes have a natural interpretation. For example, for 3D measurements of the human head, the midplane is known as the sagittal plane, and it is natural to require the coordinate axes to have the following interpretations:
  • coordinate 1: left-right
  • coordinate 2: down-up
  • coordinate 3: back-front
Figure 1 shows these coordinate axes on a human face. Further details are given in Section 4.
Figure 1: Coordinates system for human face with principal planes. (Source: Patel et al. (2023); coordinate labels have been corrected.)
To carry out axis registration it is necessary to know or estimate the midplane. There are three typical scenarios: (a) the midplane is known a priori; (b) the midplane has been estimated using expert knowledge, perhaps involving information beyond the landmarks themselves; or (c) the midplane is estimated from the landmark positions using Procrustes analysis. Scenario (c) is described in detail in Appendix A following Mardia et al. (2000) and Kent and Mardia (2001).
To carry out basis registration, it is also necessary to know or estimate the $M-1$ remaining meaningful basis vectors in $\mathcal{P}$. In this case the three possibilities become: (a) the basis is known a priori; (b) the basis is estimated using expert knowledge, perhaps involving information beyond the landmarks themselves; or (c) hybrid Procrustes analysis. For details in Procrustes analysis, see, for example Dryden and Mardia (2016). By “hybrid”, we mean that (i) the normal direction to the midplane is estimated by combining an object with its reflection and using generalized Procrustes analysis (GPA), (ii) an “average shape” is found from all the objects using GPA, (iii) the midplane for the average shape is oriented so the basis directions have natural interpretations using expert knowledge, and (iv) each data configuration is aligned to the average shape using ordinary Procrustes analysis. The details are given in Appendix A.
For the smile data application in Section 4, expert knowledge (scenario (b)) has been used to estimate the midplane with the stated interpretations for the coordinate axes as above in basis registration.
2.2 Elementary features
Suppose the size-and-shape dataset has been basis registered about a midplane $\mathcal{P}$ that passes through the origin and is normal to the first coordinate axis. Then it is possible to define various elementary features to describe the asymmetry of each data object. Paired and solo landmarks are treated separately.
Let $K_{L}$ and $K_{R}$ denote the total number of landmarks on the left or right of $\mathcal{P}$. We have $K_{L}=K_{R}=K_{P}$. Let $k_{L}$ and $k_{R}$ denote the left and right indices for a typical landmark pair for a bilateral object $X$ and consider the $M$ coordinatewise signed elementary features for the landmark pairs,
$d[(k_{L},k_{R}),1]$ $=X[k_{L},1]+X[k_{R},1]$ (2.1)
$d[(k_{L},k_{R}),m]$ $=X[k_{L},m]-X[k_{R},m],\quad m=2,\ldots,M.$ (2.2)
That is, there is one feature for each coordinate and each landmark pair. So there are in total $MK_{P}$ coordinatewise features for all paired landmarks. For this specific $\mathcal{P}$, the first coordinates for the two landmarks in one pair have opposite signs, while other coordinates have the same sign. Further, for a symmetric object, the quantities in equations (2.1) and (2.2) should all be 0. Thus, taking the sum of the first coordinates and differences for other coordinates allow us to quantify the departure from symmetry. Similarly, if $k_{S}$ is a typical solo landmark, consider the single elementary feature
$d[(k_{S})]=X[k_{S},1].$ (2.3)
There are in total $K_{S}$ elementary features for solos. Under bilateral symmetry with $\mathcal{P}$ chosen as above, the first coordinate of a solo landmark will be 0, with no restriction on other coordinates. Hence, a single feature as in (2.3) is adequate to describe the asymmetry of a solo landmark.
The coordinatewise elementary features can be collected into a signed elementary feature vector $\bm{d}=(d_{j})$, say, of length $J_{\text{basis}}=MK_{P}+K_{S}$. The elements of $\bm{d}$ will either be listed sequentially with a subscript index, i.e., $d_{j},\ j=1,\ldots,J_{\text{basis}}$, or using parentheses within square brackets, as in (2.1)-(2.3).
When only axis registration is available, the coordinate information in (2.2) does not have a well-defined interpretation. This is because the coordinates except the first one of paired landmarks would be changed after rotation within $\mathcal{P}$, hence leads to changes in values of (2.2) even though the asymmetry information does not change. Note that the equation (2.1) will not change. Further, at each landmark the coordinate information can be collected into a single number, called a landmark elementary feature which is defined per landmark,
$d^{*}[(k_{L},k_{R})]=\left\{\sum_{m=1}^{M}d^{2}[(k_{L},k_{R}),m]\right\}^{1/2},$ (2.4)
which does have a well-defined interpretation as the Euclidean distance between one landmark in a pair and the reflection of the other landmark in the pair. Rotation within $\mathcal{P}$ would not change this Euclidean distance. For a solo landmark, (2.3) is always well defined since the first coordinate of a solo remains unchanged after rotation within $\mathcal{P}$; hence it is both a coordinatewise and a landmark elementary feature. Thus, overall in the case of axis registration, a smaller set of $J_{\text{axis}}=K_{P}+K_{S}$ elementary features can be collected into a vector for further analysis. On the other hand, for $M=2$, there is no difference between axis and basis registration. Thus, equations (2.1)-(2.3) are all valid for both types of registrations in this case.
Another useful way to think about the elementary features is by using reflection. Let $X$ be a basis registered configuration, and let $X^{(\text{refl})}$ denote the reflection of $X$ about $\mathcal{P}$. Computationally, this matrix is obtained by changing the sign of the first column of $X$ and interchanging the row indices for each landmark pair. Then the elementary features of $X$ can be obtained by comparing the elements of $X$ and $X^{(\text{refl})}$. If $X$ is bilaterally symmetric about $\mathcal{P}$, then $X=X^{(\text{refl})}$ and elementary features vanish, i.e. $\bm{d}=\bm{0}$.
In many examples, such as the smile application studied later in this paper in Section 4, the direction of asymmetry, for example, left or right asymmetry, can vary between individuals and is not thought to be interested. We will focus on the absolute values of the elementary features. Define the absolute elementary feature vector $\bm{a}$ with elements
$a_{j}=|d_{j}|,\quad j=1,\ldots,J,$ (2.5)
where $J=J_{\text{basis}}$ or $J=J_{\text{axis}}$, as appropriate. In other words, only the extent of the asymmetry is interested. The elements of $\bm{a}$ can also be written in a similar way as elements of $\bm{d}$ in (2.1)-(2.3). In the rest of the paper, each configuration $X$ is reduced to an absolute elementary feature vector $\bm{a}$ for further numerical and statistical analysis.
2.3 A Simple Illustrative Example
To illustrate the construction of the vector $\bm{a}$, consider two configurations in $M=2$ dimensions with $K=4$ landmarks given by
$X_{1}=\begin{pmatrix}-1&\phantom{-}0\\ \phantom{-}0&\phantom{-}1\\ \phantom{-}1&\phantom{-}0\\ \phantom{-}0&-1\end{pmatrix},\quad X_{2}=\begin{pmatrix}-0.95&\phantom{-}0.36\\ -0.28&\phantom{-}2.11\\ \phantom{-}0.99&\phantom{-}0.54\\ -0.31&-1.37\end{pmatrix}.$ (2.6)
Suppose the landmarks consist of one landmark pair, $(k_{L},k_{R})=(1,3)$ and two solos, $k_{S}=2,4$. Hence $K_{P}=1$ and $K_{S}=2$. Further, suppose both configurations are treated as basis registered with $\mathcal{P}$ normal to the first coordinate axis, i.e. $\mathcal{P}$ is the $y$-axis.
[…]
Figure 2: The left and right figures in the top row show the original configurations of symmetric square $X_{1}$ and asymmetric square $X_{2}$ (in black) respectively. The left and right figures in the bottom row show the original objects (in black solid lines) together with their respectively reflections $X_{1}^{\text{(refl)}}$ and $X_{2}^{\text{(refl)}}$ (in red dotted lines) respectively, in order to illustrate the coordinatewise elementary feature vector.
Table 1: Example in Section 2.3. Elements of the absolute elementary feature vector $\bm{a}\in\mathbb{R}^{4}$ for $X_{2}$ with $K=4$ landmarks in $M=2$ dimensions.
Landmark Indices Coordinate Axis Feature of $\bm{a}$ value of $\bm{a}$
Pair (1,3) 1 $a[(1,3),1]=|X_{2}[1,1]+X_{2}[3,1]|$ 0.04
Pair (1,3) 2 $a[(1,3),2]=|X_{2}[1,2]-X_{2}[3,2]|$ 0.18
Solo 2 1 $a[(2)]=|X_{2}[2,1]|$ 0.28
Solo 4 1 $a[(4)]=|X_{2}[4,1]|$ 0.31
3 Hypothesis tests
Consider a dataset of $N$ basis registered configurations observed on two groups of individuals, where the first $N_{1}$ configurations belong to one group and the last $N_{2}$ configurations belong to the other group, $N=N_{1}+N_{2}$, where it is thought that the first group of configurations might be less asymmetric than the second. One way to think about this question is to construct a one-sided hypothesis test of
$\begin{split}&H_{0}:\text{ the two groups have the same distribution of asymmetry vs.}\\ &H_{1}:\text{ Group 1 is less asymmetric than Group 2.}\end{split}$ (3.1)
We will give two general approaches to construct a test statistic: (a) combine-then-compare, and (b) compare-then-combine. These two approaches will be explored in the next subsections.
To specify the null and alternative hypotheses explicitly, some notations are needed. Suppose the data take the form of $N$ registered configurations
$X_{n},\ n=1,\ldots,N.$ (3.2)
Let $\bm{a}_{n}$ denote the corresponding absolute elementary feature vector for $X_{n}$ (equation (2.5)), with elements $a_{nj},\ j=1,\ldots,J$ and $J=J_{\text{basis}}$ for the rest of this paper.
Then the hypotheses given in equation (3.1) can be reformulated as:
$\begin{split}&H_{0}:\text{ $\bm{a}_{n}\sim F$ i.i.d. for all $n$ vs.}\\ &H_{1}:\text{ $\bm{a}_{n}\sim F_{1}$ i.i.d. for $n=1,\dots,N_{1}$, $\bm{a}_{n}\sim F_{2}$ i.i.d. for $n=N_{1}+1,\dots,N$,}\end{split}$ (3.3)
where $F$, $F_{1}$, $F_{2}$ are multivariate cumulative density functions (c.d.f.s) and $F_{1}$ is stochastically smaller than $F_{2}$ (that is, $F_{1}(\bm{a})\geq F_{2}(\bm{a})$ for all $\bm{a}$ with at least one strict inequality), see, for example Stoyan (1983). In other words, the distribution of $F_{2}$ is shifted towards right comparing with $F_{1}$, which indicates more asymmetries.
We concentrate in this section on t-test and Mann-Whitney U test. In Section 4, we further discuss some alternative univariate tests as well as multivariate tests in relation to the smile data.
3.1 Testing strategy: combine-then-compare
Each element $a_{j}$ of $\bm{a}$ (equation (2.5)) is nonnegative and a larger value of $a_{j}$ indicates greater asymmetry. Hence it is natural to reduce the $J$-dimensional vector $\bm{a}$ to a single number by combining the elements of $\bm{a}$ into a composite asymmetry score
$u=\phi(\bm{a}),$ (3.4)
say. Here the function $\phi(\bm{a})$ is assumed to be monotonically increasing in the sense that it is monotonically increasing in each component $a_{j}$ when the other elements are held fixed.
Let $\psi(a)$ be a monotonically nonegative increasing function of a scalar argument $a\geq 0$. Then one way to define $\phi(\bm{a})$ is as an additive composite asymmetry score
$\phi_{\psi,\bm{w}}(\bm{a})=\sum_{j=1}^{J}w_{j}\psi(a_{j}),$ (3.5)
where the $w_{j}\geq 0$ are pre-assigned weights. Some choices for weights include the following:
  • (a)
    equal weights $w_{j}=1$,
  • (b)
    greater weights near the midplane,
  • (c)
    greater weights away from the midline, or
  • (d)
    different weights for pairs and solos.
For example, in cases (b) and (c) a weight might depend on $|X[k_{R},1]-X[k_{L},1]|$ for a landmark pair, with solos treated separately. The choices of weights in general depend on prior belief of the extent of asymmetry information carried by each landmark, for example, if it is known a priori that landmarks far away from $\mathcal{P}$ can carry important asymmetry information than landmarks near $\mathcal{P}$, then more weights should be assigned to the farther landmarks. For smile example, this information is available so greater weights will be given near the midplane, see Section 4.
The choice of function $\psi$ governs the relative influence of different magnitudes of an elementary feature. For example, the quadratic function $\psi(a)=a^{2}$ is more sensitive to outlying features than the linear function $\psi(a)=a$.
We write below explicitly for two such scores with equal weights and give them names $L_{1}$ and $L_{2}$ statistics respectively.
[…]
In the above equations, we use the two expressions for elements in $\bm{d}$ and $\bm{a}$ introduced in Section 2.2.
Bock and Bowman (2006) have proposed an asymmetry score with weights $w_{j}=1$ for landmark pairs and weights $w_{j}=2$ for solos and $\psi(a_{j})=a_{j}^{2}$ in equation (3.5), namely,
$\sum_{(k_{L},k_{R})}\sum_{m=1}^{M}d[(k_{L},k_{R}),m]^{2}+2\sum_{k_{S}}d[(k_{S})]^{2}.$ (3.8)
For our work, we have selected an equal weights for $\phi_{L_{1}}$ to have some similarity to the asymmetry score of Patel et al. (2023); their score is given by
$\phi_{L_{1}}^{*}(\bm{a})=\sum_{(k_{L},k_{R})}d^{*}[(k_{L},k_{R})]+\sum_{k_{S}}|d[(k_{S})]|,$ (3.9)
with weights equal to 1 for both pairs and solos. The $d^{*}[(k_{L},k_{R})]$ is defined in equation (2.4). Also, we can define a similar score as follows using $L_{2}$ distances
$\phi_{L_{2}}^{*}(\bm{a})=\sum_{(k_{L},k_{R})}d^{*}[(k_{L},k_{R})]^{2}+\sum_{k_{S}}|d[(k_{S})]|^{2}.$ (3.10)
It can be shown using (2.4) that this equation (3.10) reduces to (3.7). That is, we have
$\phi^{*}_{L_{2}}(\bm{a})=\phi_{L_{2}}(\bm{a}).$
Note that, we can generalize $\phi^{*}(\bm{a})$ in the same way as $\phi(\bm{a})$ with weights.
Example in Section 2.3 continued. Recall the asymmetric square $X_{2}$ defined in equation (2.6). The configuration is shown in Figure 2. The landmark elementary feature, $d^{*}[(1,3)]$, defined in equation (2.4) is computed and reported in Table 2. The composite asymmetry scores $\phi_{L_{1}}(\bm{a})$ (3.6), scaled $\phi^{*}_{L_{1}}(\bm{a})$ (3.9) (divided by the number of landmarks) and $\phi_{L_{2}}(\bm{a})$ (3.7) are computed on $X_{2}$ for illustration. The results are shown in Table 2. Note that the values of
$d^{*}[(1,3)]=\sqrt{(X_{2}[1,1]+X_{2}[3,1])^{2}+(X_{2}[1,2]-X_{2}[3,2])^{2}}=0.18$
and $a[(2)]=0.28$, $a[(4)]=0.31$ (Table 1) are less than 1, so taking squares of them in $\phi_{L_{2}}$ leads to smaller values. Note that here we have scaled $\phi^{*}_{L_{1}}$ by a factor of half, so unscaled $\phi^{*}_{L_{1}}=0.78$ which is similar to the value of $\phi_{L_{1}}=0.81$.
$d^{*}[(1,3)]$ (2.4) $\phi_{L_{1}}(\bm{a})$ (3.6) $\phi^{*}_{L_{1}}(\bm{a})$ (3.9) $\phi_{L_{2}}(\bm{a})$ (3.7)
$X_{2}$ 0.18 0.81 0.39 0.21
Table 2: The two composite asymmetry scores defined in equation (3.5) and equation (3.9) are computed on the configuration $X_{2}$.
Once a composite asymmetry score $\phi(\bm{a})$ has been chosen, let
$u_{n}=\phi(\bm{a}_{n}),\quad n=1,\ldots,N,$ (3.11)
denote the composite asymmetry scores for the data in (3.2).
Let
$\mu_{g}=E\{U_{g}\},\ \sigma^{2}_{g}=\text{var}\{U_{g}\}$ (3.12)
denote the expectation and variance for a random variable $U_{g}\sim F_{g}$, where $g=1,2$ labels one of the two groups. A simpler version of hypotheses in (3.1) is
$H_{0}:\mu_{1}=\mu_{2}\text{ vs. }H_{1}:\mu_{1}<\mu_{2}$ (3.13)
There are several test statistics that can be used for testing hypotheses given in (3.1) and (3.13), including the following:
  • (a)
    the standard two-sample t-test, which assumes a common variance for the two groups ($\sigma_{1}^{2}=\sigma_{2}^{2}$);
  • (b)
    the Mann-Whitney U test, also known as the Wilcoxon rank sum test (Mann and Whitney, 1947).
The Mann-Whitney U test is nonparametric so that the test results are invariant under monotonic transformations of the $u_{n}$. In case (a), the significance of the test statistic under $H_{0}$ can be computed either analytically (assuming normality) or using a bootstrap approximation. In case (b) the significance can be derived from combinatorial arguments. Mann-Whitney U test is considered where the normality assumption on composite asymmetry scores $u_{n}=\phi(\bm{a}_{n})$ may fail to hold.
An advantage of the combine-then-compare approach is that it enables us to accumulate evidence from different features. If there is a small amount of asymmetry on many different elementary features, then the composite score will have a large value. Further, this approach enables the two groups to be compared visually with stem and leaf plots and in particular, any overlap between the two groups can be easily assessed. Applications of this approach are in Section 4
3.2 Testing strategy: compare-then-combine
In this approach, separate test statistics are constructed for each feature in the elementary feature vectors. The most extreme of these separate test statistics is then used as an overall test statistic. Before we give the details, the hypotheses in equation (3.1) need to be formulated in the union intersection test style (as explained in Section 1) as following:
$H_{0}=\bigcap_{j=1}^{J}H_{0}^{j}\text{ vs }H_{1}=\bigcup_{j=1}^{J}H_{1}^{j},$ (3.14)
where $H_{0}^{j}$: the mean of $j$th unsigned elementary feature between the two groups is the same, whereas $H_{1}^{j}$ is that the mean of group 2 is larger than mean of group 1, for $j=1,\dots,J$.
For each choice $j$ of an unsigned elementary feature, construct a statistic to compare the two groups, for example, a two-sample t-statistic or a Mann-Whitney U statistic. Denote the resulting statistics by $v_{j},\ j=1,\ldots,J$. For example, if a t-statistic is used, then
$v_{j}=(\bar{a}_{j}^{(1)}-\bar{a}_{j}^{(2)})/s_{j}.$ (3.15)
where $\bar{a}_{j}^{(g)},\ g=1,2$ are the sample means of the $j$th unsigned elementary feature in the two groups and $s^{2}_{j}$ is the pooled within-group variance.
An overall UIT statistic to test $H_{0}$ vs $H_{1}$ in equation (3.14) can be defined by taking the maximum of these separate statistics (Boyett and Shuster, 1977),
$V=\max_{j=1,\ldots,J}v_{j}.$ (3.16)
The critical points of $V$ under the null hypothesis can be computed by the bootstrap. Note that the landmarks are correlated with each other, so the $d_{j}$ are not independent for $j=1,\ldots,J$ here, so we cannot derive the distribution of $V$ theoretically. Hence, bootstrap is used for estimating the critical points of $V$.
If there exists $v_{j}$ which exceeds the critical values of $V$, then we not only know that $H_{0}$ is rejected, but also know that the landmark corresponding to such $v_{j}$ is also important.
Our main aim is somewhat different from Boyett and Shuster (1977): we aim at selecting important landmarks while Boyett and Shuster (1977), attempt to estimate a $p$-value. The bootstrap procedure used by us is also different from their who have used:
  1. 1.
    A set $E$ containing all permutations of the dataset is constructed.
  2. 2.
    Sample with replacement from this set $E$, i.e. sample permutations of the data. The permuted samples constitute the bootstrap resampled dataset.
  3. 3.
    Test statistic $V$ is computed on the resampled dataset.
On the other hand, we directly sample from the whole dataset as following:
  1. 1.
    Sample $N$ samples with replacement from $\{\bm{a}_{n}\}_{n=1}^{N}$.
  2. 2.
    Compute $v_{j}$ and $V$ on resampled dataset.
An advantage of this compare-then-combine approach versus combine-and-compare is that if the null hypothesis is rejected, then simultaneous confidence intervals can be constructed to identify those features on which the two groups significantly differ. The selection procedure in spirit is similar to Tukey’s method (Tukey, 1949) in ANOVA though in ANOVA a separate test statistic is used for each pairwise difference between main effects. Here, a separate test statistic for each $j$ is more appropriate since all features in unsigned elementary feature vector $\bm{a}$ have different biology. An illustration is given below for the smile data.
4 Analyses of the Smile Data
We now give smile data taken from Patel et al. (2023) which have been pre-registered. The data is collected using the Di4D system. When a subject’s face is placed in front of the camera, the $x$-axis is positioned horizontally on the face (positive direction is left), the $y$-axis is placed vertically on the face (positive direction is upward), whereas the $z$-axis is positioned in the direction inwards and outwards of the face (positive direction is outward), see Figure 1.
The coordinates given in Section 2.1 can be related to these coordinate axes as follows:
coordinate 1: left-right $x$
coordinate 2: down-up $y$
coordinate 3: back-front $z$.
Our smile data contains $N=25$ subjects, of which $N_{1}=12$ are control subjects and $N_{2}=13$ are cleft subjects. There are $K=24$ landmarks on the lip periphery. Figure 3 shows the lip configuration in the $x$-$y$ plane. There are $K_{P}=11$ landmark pairs and $K_{S}=2$ solo landmarks. The landmark indices are summarized in Table 3. The $M=3$ dimensional coordinates of these landmarks have been extracted at three frames: first (closed lip), middle (middle of the smile) and last frame (maximum open lip smile). Let $g=1,2$ denote the control and cleft groups, respectively.
Figure 3: Landmark indices for $x$-$y$ coordinates on the lip periphery of a control subject at first frame.
Table 3: Indices of landmark pairs and solos in Figure 3 on the lip periphery.
Landmark notation Indices in Figure 3
pair $(k_{L},k_{R})$ (1,13), (2,12), (3, 11), (4,10), (5,9), (6,8),
(20,18), (21,17), (22,16), (23,15), (24,14)
$k_{S}$ 7, 19
In the rest of this section several test statistics will be presented in two parts; combine-then-compare approach and compare-then-combine approach.
Combine-then-compare approach using (3.9). Patel et al. (2023) carried out standard two-sided two-sample t-tests using the statistic (3.9), with conventional t-tables used to judge significance. Here we report the results of the one-sided versions of these t-tests since based on hypotheses in (3.1), one-sided tests are more appropriate. The results of our t-tests are given in Table 4. We reach the same conclusions as in Patel et al. (2023) for their two-sided tests, namely, $H_{0}$ is rejected at all three frames. Further, the one-sided t-test indicates that there is more asymmetry for cleft lip subjects versus controls. Note that here the $p$-value is the smallest at the first frame, which suggests that the two groups are the most different at the start.
Table 4: Mean, variance, t-values and $p$-values for $\phi^{*}_{L_{1}}(\bm{a})$ (equation (3.9)) for cleft and control subjects at first, middle and last frames (Patel et al. (2023)).
Cleft Control One-sided test
mean sd mean sd $t$-values $p$-values
First 1.78 0.46 1.12 0.41 3.86 0.0004(***)
Middle 1.96 0.70 1.46 0.47 2.14 0.02(*)
Last 2.14 0.80 1.55 0.43 2.34 0.01(**)
(*) = significant at the $5\%$ significance level. (**) = significant at the $1\%$ significance level. (***) = significant at the $0.1\%$ significance level.
Combine-then-compare approach using (3.6) and (3.7). We use the scores $\phi_{L_{1}}(\bm{a})$ and $\phi_{L_{2}}(\bm{a})$ defined in equations (3.6) and (3.7) respectively. Instead of considering asymmetries with respect to each landmark as in (3.9), we view all features in $\bm{a}$ as a whole and take the $L_{1}$ and $L_{2}$ norms. For the subjects which are more symmetric, their corresponding $\bm{a}_{n}$ should be close to the origin. Hence, the $L_{1}$ and $L_{2}$ distances between $\bm{a}_{n}$ with $\bm{0}\in\mathbb{R}^{J}$ should be smaller for control subjects. Several one-sided two-sample Mann-Whitney U tests have been performed and the results are shown in Table 5. For the last two rows in the table, the weighted $\phi_{L_{1}}(\bm{a})$ and weighted $\phi_{L_{2}}(\bm{a})$ are used and $w_{j}$ are chosen in an adaptive way as the following:
  1. 1.
    Compute the sample mean shape on the augmented data, which contains basis registered original configurations and their reflections through midplane. Since the data has been pre-registered, the sample mean shape is simply the arithmetic mean for each landmark.
  2. 2.
    The Euclidean distance within each landmark pair for this mean shape is computed and its reciprocal is used as the weight. The weights are the same among these subjects from both groups. The weights for solo landmarks are the unit length based on the scale of data.
Different landmarks carry different amount of asymmetry information. Those landmarks near the midplane $\mathcal{P}$ may carry the most significant information on asymmetry according to the expert’s knowledge. Thus, we give higher weights to landmarks near $\mathcal{P}$ and smaller weights on landmarks faraway from the central.
Table 5: $p$-values from one-sided two-sample Mann-Whitney U tests using $\phi_{L_{1}}(\bm{a})$ (3.6) and $\phi_{L_{2}}(\bm{a})$ (3.7).
Composite asymmetry score First frame Middle frame Last frame
$\phi_{L_{1}}(\bm{a})$ $9.69\times 10^{-5}$(***) 0.026(*) 0.009(**)
$\phi_{L_{2}}(\bm{a})$ 0.0002(***) 0.011(*) 0.013(*)
weighted $\phi_{L_{1}}(\bm{a})$ 0.002(**) 0.008(**) 0.002(**)
weighted $\phi_{L_{2}}(\bm{a})$ 0.013(*) 0.015(*) 0.005(**)
(*) = significant at the $5\%$ significance level. (**) = significant at the $1\%$ significance level. (***) = significant at the $0.1\%$ significance level.
It can be seen from Table that 5 $H_{0}$ is rejected in all cases. In summary, there are significant differences between cleft and control subjects at all frames, with control subjects less asymmetric than cleft subjects overall.
Since the test results indicate significant differences between the two groups, we would like to visualize the separation between the distributions of composite asymmetry scores of the two groups. Hence, dot plots similar to the stem leaf plot are created and given in Figures 4, 5, 6, 7 and 8. In these dot plots, composite asymmetry scores for both groups are plotted together at each frame.
Figure 4 shows the dot plots for $\phi^{*}_{L_{1}}(\bm{a})$ defined in equation (3.9), whereas Figure 5 and 6 display the dot plots of $\phi_{L_{1}}(\bm{a})$ and $\phi_{L_{2}}(\bm{a})$ (equations (3.6) and (3.7)) respectively. According to these dot plots, substantial overlaps between cleft and control subjects can be seen. Moreover, the three composite asymmetry scores used in this section all have similar performances, as the separation between the two groups at each frame revealed by the dot plots are all similar.
i First frame
ii Middle frame
iii Last frame
Figure 4: Dot plots of $\phi^{*}_{L_{1}}(\bm{a})$ defined in equation (3.9) at the three frames. The top row of the graph is for cleft lip subjects whereas the bottom row is for controls.
i First frame
ii Middle frame
iii Last frame
Figure 5: Dot plots of $\phi_{L_{1}}(\bm{a})$ (3.6) at the three frames. The top row of the graph is for cleft lip subjects whereas the bottom row is for control subjects.
i First frame
ii Middle frame
iii Last frame
Figure 6: Dot plots of $\phi_{L_{2}}(\bm{a})$ (3.7) at the three frames. The top row of the graph is for cleft lip subjects whereas the bottom row is for control subjects.
Figure 7 and 8 show the dot plots for weighted $\phi_{L_{1}}(\bm{a})$ and weighted $\phi_{L_{2}}(\bm{a})$ respectively. According to these figures, these two composite scores push outliers farther away, i.e., they are sensitive to outliers, especially for weighted $\phi_{L_{2}}(\bm{a})$. So the weighted $\phi_{L_{1}}(\bm{a})$ and weighted $\phi_{L_{2}}(\bm{a})$ emphasize which are more extreme in cleft lip group.
i First frame
ii Middle frame
iii Last frame
Figure 7: Dot plots of weighted $\phi_{L_{1}}(\bm{a})$ at the three frames. The top row of the graph is for cleft lip subjects whereas the bottom row is for control subjects.
i First frame
ii Middle frame
iii Last frame
Figure 8: Dot plots of weighted $\phi_{L_{2}}(\bm{a})$ at the three frames. The top row of the graph is for cleft lip subjects whereas the bottom row is for control subjects.
Compare-then-combine approach using (3.4). The method described in Section 3.2 is used to select the landmarks which possess significant asymmetry information. $v_{j}$ is selected to be the two-sample t-statistics (equation (3.15)). $V$ defined in (3.16) is computed and its critical value at $5\%$ significance level is determined via bootstrap. The total number of bootstrap iterations used was 10000. Boyett and Shuster (1977) suggest 17000 iterations to make sure the significance level is $1\%$. However, the results of using 10000 and 17000 iterations turned out to be the same for our smile data. The central pair, landmarks 6 and 8, correspond to $t$-values which exceed these critical points at all three frames. Thus, these two landmarks have the most significant information of asymmetry. Further, the simultaneous confidence interval related to landmarks 6 and 8 excludes 0, so that $H_{0}$ is rejected when use these landmarks.
Further, most of the extreme $p$-values are obtained from Mann-Whitney U tests related to landmarks 6, 8 and two solo landmarks 7 and 19. These tests are performed using $\phi_{L_{1}}$ and $\phi_{L2}$ computed on various subsets of landmarks. This suggests that the two groups are differed at most when using these four landmarks. So these landmarks can carry important information on asymmetries. This finding matches the expert’s opinion. Figure 9 shows a photo of a cleft patient, where the central pair (6, 8) and solo landmarks, 7 and 19, are marked on the photo and the landmarks are same as in Figure 3. Roughly, it is a ‘Y’-shape if we join the landmarks 6, 7, 8 and then 7 with 19 by line segments. From medical point of view, this particular shape is known a priori to be the most important for the asymmetry assessment.
Figure 9: Photo of a cleft patient with central pair landmarks and solo landmarks. This figure is reproduced from photo taken from the website https://www.nhs.uk/conditions/cleft-lip-and-palate/.
We also applied the Welch version of the $t$-test (Welch, 1947), which accommodates different variances for the two groups. In this example, the results of the Welch $t$-test is found to be similar to the standard $t$-test (Table 4). A possible reason is that our sample size is small, so it results into similar outcome. Further, the multivariate Hotelling’s $T^{2}$ test on unsigned elementary feature vector $\bm{a}_{n}$, $n=1,\ldots,N$, is an alternative approach to the UIT. In this example, the test results fail to improve the outcomes of the univariate tests. It could be since a Hotelling’s $T^{2}$ test treats the features separately so it does not gain any power if the features point in the same direction; one-sided Hotelling’s $T^{2}$ could be appropriate approach, which we will pursue in future. In this example, the sample size is less than the dimension of $\bm{a}_{n}$, i.e. $N<J$, so the covariance matrix is singular, which requires some modifications and will be dealt in future.
We end this section by summarizing the main conclusions for this smile data:
  • There are statistically significant differences between the two groups at all frames. So overall, control subjects are less asymmetric than the cleft lip subjects.
  • Substantial overlaps can be found on dot plots in Figure 4, 5 and 6. This matches the expert’s expectation that some cleft lip subjects should be fairly close to normal subjects after surgeries.
  • The landmarks 6, 7, 8 and 19 (the ‘Y’-shape) are the most important in assessing asymmetries.
5 Discussion
We have presented our work for pre-registered data, but it can be extended when the landmark data is not pre-registered as pointed out in Section 4. Further, these revised measures of asymmetry can be applied and analyzed in a straightforward way. Further, we have used only size-and-shape framework but it can be extended easily to include for example, similarity shape.
We have used absolute values ($a_{j}$) on coordinate elementary features as in equation (2.5) which is of main interest if one is looking for the extent of asymmetry as in our case for the smile data. However, there maybe some other applications where raw $d_{j}$ may be important than $a_{j}$; our work easily applies to that situation.
In this paper we have concentrated on smile example with three fixed frames. Our future statistical work will be extended to dynamic shape analysis which will allow to study trajectories of the smile. Some initial work for a single group of a smile data has been introduced in Mardia et al. (2018) and Bookstein et al. (2024). We have concentrated on the cleft lip subjects but there are other medical applications such as in facial reconstruction related to trauma surgery where the lips form significant area for surgery.
We have used univariate methods but we can use Hotelling’s $T^{2}$ in place of t-statistics, for example, though it requires some modifications. An one-sided alternative seems natural according to equation (3.1) and in future we would consider one-sided multivariate tests. Nevertheless, it requires more care to set up than in the one-dimensional case. Further, the scenario of sample size is less than the dimension of the data vector has also to be taken into account.
Acknowledgements
The authors would like to thank the School of Mathematics, University of Leeds for supporting the smile project. Kanti Mardia acknowledges the Leverhulme Trust for the Emeritus Fellowship.
Appendix A Estimation of midplane using Procrustes analysis
Let $\bm{n}\in\mathbb{R}^{M}$ denote the unit normal vector of an arbitrary plane $\mathcal{P}^{\prime}$. Let $X_{n}\in\mathbb{R}^{K\times M}$, $n=1,\dots,N$ denote the observed configurations. The reflection of $X_{n}$ about $\mathcal{P}^{\prime}$, $X^{\text{(refl)}}_{n}$, is given as
$X^{\text{(refl)}}_{n}=X_{n}H,$
where $H=I_{M}-2\bm{n}\bm{n}^{T}$ is the Householder matrix and $I_{M}$ is the $M\times M$ identity matrix. Then the estimation process used in basis registration is given as the following:
  1. 1.
    Form an augmented dataset: $\{X_{n},X^{\text{(refl)}}_{n}\}_{n=1}^{N}$.
  2. 2.
    Apply GPA on this augmented dataset. Denote the fitted configurations as $X_{n}^{(\text{GPA})}$ for $n=1,\dots,N$.
  3. 3.
    Compute the Procrustes mean shape. This mean shape will be bilaterally symmetric with midplane $\mathcal{P}^{\prime}$.
  4. 4.
    Rotate the mean shape outside and within $\mathcal{P}^{\prime}$ until the new midplane and the basis are meaningful.
  5. 5.
    Carry out ordinary Procrustes analysis between each $X_{n}^{(\text{GPA})}$ and the rotated mean shape.
When axis registration is considered, we only rotate the sample mean shape outside midplane. Further, the last step, step 5, is ignored. Further details are given for this procedure in Mardia et al. (2000).

### v2

@@
 Abstract
-Many biological objects possess bilateral symmetry about a midline or midplane, up to a “noise” term.
-This paper uses landmark-based methods to measure departures from bilateral symmetry, especially for the two-group problem where one group is more asymmetric than the other.
-In this paper, we formulate our work in the framework of size-and-shape analysis including registration via rigid body motion.
-Our starting point is a vector of elementary asymmetry features defined at the individual landmark coordinates for each object.
-We introduce two approaches for testing.
-In the first, the elementary features are combined into a scalar composite asymmetry measure for each object.
-Then standard univariate tests can be used to compare the two groups.
-In the second approach, a univariate test statistic is constructed for each elementary feature.
-The maximum of these statistics lead to an overall test statistic to compare the two groups and we then provide a technique to extract the important features from the landmark data.
-Our methodology is illustrated on a pre-registered smile dataset collected to assess the success of cleft lip surgery on human subjects.
-The asymmetry in a group of cleft lip subjects is compared to a group of normal subjects, and statistically significant differences have been found by univariate tests in the first approach.
-Further, our feature extraction method leads to an anatomically plausible set of landmarks for medical applications.
-Keywords— asymmetry scores, cleft lip disfigurement, facial trauma surgery, feature selection, registration, shape analysis, size-and-shape, smile.
+Many biological objects possess bilateral symmetry with some noise about a midline or midplane.
+This paper uses landmark-based methods to measure departures from bilateral symmetry, especially for paired data and two-group problems where one group may be more symmetric than the other.
+We formulate our work under size-and-shape analysis framework including pre-registration under rigid body motion.
+We then construct a vector of elementary asymmetry features defined per landmark coordinate per object.
+Next, we introduce two approaches for comparing bilateral symmetry between the two groups or paired data: (1) the elementary features are combined into a scalar asymmetry measure for each object, then standard univariate tests are used to test for asymmetries; (2) a univariate test statistic is constructed for each feature, then the maximum of these statistics leads to an overall test statistic and important features can be extracted from the data.
+Our methodology is illustrated on two registered datasets collected to assess the success of two kinds of surgeries: (1) cleft lip surgery data, where the symmetry is compared between a group of cleft lip subjects and normal subjects; (2) orthognathic surgery data, where the patients’ symmetry after the orthognathic surgery are compared to that before surgery.
+Keywords: asymmetry measures, cleft lip surgery, orthogonathic surgery, registration, shape analysis, size-and-shape analysis of smile.
 1 Introduction
-Bilateral symmetry (left-right symmetry) is a key property in many biological settings.
-An object (treated as a set of unlabelled points known as landmarks) is said to be bilaterally symmetric if its reflection through some midline or midplane $\mathcal{P}$ is exactly the same as the original object.
+Bilateral symmetry, also called left-right symmetry, is a key property in many biological settings.
+An object is said to be bilaterally symmetric if its reflection through some midline (for objects in two dimensions) or midplane (for objects in three dimensions) passing through the center of the object is exactly the same as the original object.
 In practice, exact bilateral symmetry seldom holds and it is of interest to study the extent of any asymmetry.
-Research about the symmetries of objects in the real world has a long history (see, for example, Weyl, 1952).
-Of particular interest is the hypothesis of bilateral symmetry, especially in biology (Palmer and Strobeck, 1986).
+Research on the symmetries of objects in the real world has a long history (see, for example, Weyl, 1952).
 Statistical methods to test asymmetry were first formalized in Mardia et al.
@@
 The cases $M=2$ and $M=3$ are the most important in practice.
-Call an object bilateral if its landmarks can be divided into two categories: pairs and solos, where paired landmarks locate on both sides of $\mathcal{P}$, while solos are unpaired.
-For a bilaterally symmetric object the two landmarks in a pair are equally spaced about the midplane and the solos lie on the midplane.
-If there are $K_{P}$ pairs of landmarks and $K_{S}$ solo landmarks, then the total number of landmarks is $K=2K_{P}+K_{S}$.
-If the dataset has not been pre-registered, then registration is required.
-The details of registration are given in Section 2.1.
-The asymmetry information in an object can be represented as a vector of coordinatewise elementary asymmetry features, $\bm{d}$, say, defined in more details in Section 2.2.
-Increasing the magnitude of any element of this vector indicates greater asymmetry.
+We call an object bilateral if its landmarks can be divided into two categories: pairs and solos.
+An object is bilaterally symmetric about an M1M-1-dimensional midplane $\mathcal{P}$ if all the solos lie on the midplane, and the two landmarks in each pair are equally spaced about $\mathcal{P}$, meaning that their midpoint lies on $\mathcal{P}$ and the vectors to each from their midpoint are orthogonal to $\mathcal{P}$.
+Let there be $K_{P}$ pairs and $K_{S}$ solos; then $K=2K_{P}+K_{S}$.
+After we have determined the $M$ orthogonal axes for the object, the intersection point between the midplane and the first coordinate axis can be computed.
+Under translation and rotation it is ensured that the midplane $\mathcal{P}$ passes through the origin and is orthogonal to the first coordinate axis.
+All objects in the dataset should be registered so that all of them share the same midplane $\mathcal{P}$, hence, further analyses are available.
+If the dataset has not been pre-registered before the analyses, then registration is required; the details are given in Section 2.1.
+This paper is motivated by two medical questions related to the level of asymmetry in a person’s smile.
+One is related to cleft lip surgery which is performed so that form and function can be restored to particularly lip symmetry at rest and during movement.
+The second is related to orthognathic surgery, which is performed to correct jaw discrepancies in order to restore the form and function but the surgery can influence facial and lip symmetry.
+We have been provided with appropriate data in each case.
+We will refer the cleft lip data as Smile Data 1, and the orthognathic surgery data as Smile Data 2.
+In both datasets, we have a set of landmarks on the lip area as appropriate to each case; Smile Data 1 contains 24 landmarks over the lip while Smile Data 2 contains 16 landmarks over the lip and nose.
+More details of the data are provided in Section 5 and Section 6 respectively.
+In each case, the subjects were asked to smile from rest (close lip) to maximum open lip smile and a four-dimensional (three spatial dimensions plus time) movie capture was made of the face using Di4D motion capture system (https://di4d.com).
+A superimposed finite element mesh combined with optical flow techniques were used to follow the landmarks through the frames of the capture, taken at one-second intervals, of the capture.
+Note that these include both anatomical landmarks, e.g.
+the corners of the mouth, and pseudo landmarks, e.g.
+a midpoint of the upper right lip defined at first frame and then followed through successive frames of the capture by optical flow.
+Smile Data 1 contains two groups: normal subjects and subjects who had previous cleft lip surgery.
+Smile Data 2 includes a single group of subjects before and after orthognathic surgery, so it is a paired dataset.
+The aim is to assess the success of these two different surgeries, that is, are the cleft lip subjects more symmetric after the surgery, and has orthognathic surgery had an impact on symmetry.
+Both datasets are pre-registered in the same way and we have three fixed frames (the beginning, middle and end of the smile) for each subject.
+The asymmetry information in an object is represented as a vector of coordinatewise elementary asymmetry features, $\bm{d}$, say, defined in Section 2.2.
+Increasing the magnitude of any element in $\bm{d}$ indicates greater asymmetry.
 The elementary feature vector is also the starting point of Bock and Bowman (2006) and Patel et al.
-(2023) but we define its elements explicitly in Section 2.2.
+(2023), and we define its elements explicitly in Section 2.2.
 Given two groups of objects, a natural question is whether one group is more asymmetric than the other.
-Section 3 looks at two testing strategies.
-In the first strategy the elements of the feature vector $\bm{d}$ are combined into a single number called a composite score.
-Then a univariate test such as the two-sample t-test or the Mann-Whitney U test can be used to compare the two groups.
-The use of a univariate summary statistic for each object facilitates the comparison between objects and between groups of objects.
-The mathematical details are given in Section 3.1, our composite score extends the asymmetry measures used by Bock and Bowman (2006) and Patel et al.
-(2023).
-In the second strategy, the two groups are compared using each of the elements in $\bm{d}$, yielding a collection of test statistics.
-The maximum of these test statistics gives an overall test statistic that can be used to test for a difference between the two groups.
-This approach is similar to the union-intersection test (UIT) in multivariate analysis, see, for example, Mardia et al.
-(2024).
-In particular, when the null hypothesis is rejected, it is possible to investigate which elementary features are responsible.
-The mathematical details are given in Section 3.2.
-The motivation for the paper is given by a set of smile data on human subjects.
-Each subject was asked to smile, and a four-dimensional movie was made of the face (three spatial dimensions plus time using Di4D motion capture system).
-A set of landmarks was identified around the lips and followed through the frames of the movie.
-The subjects come from two groups: normal subjects and subjects who have had cleft lip surgery.
-The objective is to assess how successful the surgery has been; that is, are the cleft lip subjects any more asymmetric than normal people.
-The details of the smile data are provided in Section 4.
-The data is pre-registered and the objects at three fixed frames (the beginning, middle and end of the smile) are analyzed in Section 4 following Patel et al.
-(2023).
-We end the paper with some discussion in Section 5.
-The procedures for estimating the midline or midplane are given in Appendix A.
+Section 3 looks at two testing strategies which we name combine-then-compare and compare-then-combine respectively.
+Our methodology involves simultaneous comparisons, hence, meta-analysis is considered.
+The literature on meta-analysis as required in the paper is scattered over several papers, and in Section 4, we give an overview of meta-analysis as required for our work.
+The meta-analysis is carried out in Section 5 and 6 respectively for the two datasets.
+The paper concludes with a discussion in Section 7.
 2 Describing asymmetry
 2.1 Registration
-We will introduce the following notations: let $X$, with elements $X[k,m]$, be a $K\times M$ configuration matrix giving the positions of $K$ landmarks in $M$ dimensions for a single object or subject or individual.
-We write as $X[k,\ ]\in\mathbb{R}^{M}$ for the coordinates of $k$th landmark, $k=1,\dots,K$.
-The most important choice for $M$ is $M=3$, and this case is emphasized in the presentation, but the mathematical theory is valid for any $M\geq 1$.
-In general, upper case letters are used for matrices and bold lower case letters for vectors.
-In addition the letters $K,M,J$ are reserved for the number of landmarks, number of dimensions, and number of elementary features (Section 2.2), respectively, with typical indices given by the lower case letters $k,m,j$.
-In general, each data object $X$ will be viewed as a noisy version of a bilaterally symmetric configuration.
-Rigid body transformations If $X$ is a configuration, then a rigid body motion takes $X$ to $\bm{1}_{M}\bm{c}^{T}+XR=X^{*}$, say, where $\bm{c}$ is an $M$-dimensional translation vector and $R$ is an $M\times M$ rotation matrix.
-For the purposes of this paper, we say that $X$ and $X^{*}$ have the same size-and-shape following the standard terminology in shape analysis (see, for example, Dryden and Mardia (2016)).
-That is, the size-and-shape of $X$ is the equivalence class of configurations under the group of rigid body motions.
+[…]
+Rigid body transformations If $X$ is a configuration, then a rigid body motion takes $X$ to $X^{*}=\bm{1}_{M}\bm{c}^{T}+XR$, say, where $\bm{c}\in\mathbb{R}^{M}$ is a translation vector and $R$ is an $M\times M$ rotation matrix.
+For the purposes of this paper, we say that $X$ and $X^{*}$ have the same size-and-shape, as the size-and-shape of $X$ is the equivalence class of configurations under the group of rigid body motions (see, for example, Dryden and Mardia (2016)).
 The asymmetry information carried by the size-and-shape is of interest and such information is unaffected under rigid body motion.
@@
 Under a rigid body motion, $\bm{n}$ is transformed to $R^{T}\bm{n}$ and $b$ is transformed to $b+\bm{n}^{T}R\bm{c}$.
-Two types of Registration For every bilateral data object $X$, suppose there is an associated midplane $\mathcal{P}$.
-Ways in which $\mathcal{P}$ might be determined are disussed below.
-For data analysis, it is helpful to “register” $X$ using a rigid body motion so that after transformation $\bm{n}=\bm{e}_{1}$ and $b=0$, i.e., the midplane passes through the origin and the normal direction is given by the first coordinate axis, while still keeps the shape of $X$ unchanged.
-We will distinguish two types of registration:
-  1. 1.
-    Axis registration
-  2. 2.
+Two types of Registration We distinguish two types of registrations
+  1. (a)
+    Axis registration.
+  2. (b)
     Basis registration.
-Axis registration takes the form described in the previous paragraph.
-In terms of the human head, after registration, the left-right axis lies along the first coordinate axis.
-However, the position of the down-up and back-front axes are not clearly distinguished from one another.
-It is possible to tilt the head forwards or backwards while still preserving axis registration.
-In some settings it is possible to carry out a stronger version of registration, we will call basis registration, in which all the coordinate axes have a natural interpretation.
-For example, for 3D measurements of the human head, the midplane is known as the sagittal plane, and it is natural to require the coordinate axes to have the following interpretations:
-  • coordinate 1: left-right
-  • coordinate 2: down-up
-  • coordinate 3: back-front
-Figure 1 shows these coordinate axes on a human face. Further details are given in Section 4.
-Figure 1: Coordinates system for human face with principal planes. (Source: Patel et al. (2023); coordinate labels have been corrected.)
-To carry out axis registration it is necessary to know or estimate the midplane.
-There are three typical scenarios: (a) the midplane is known a priori; (b) the midplane has been estimated using expert knowledge, perhaps involving information beyond the landmarks themselves; or (c) the midplane is estimated from the landmark positions using Procrustes analysis.
-Scenario (c) is described in detail in Appendix A following Mardia et al.
-(2000) and Kent and Mardia (2001).
-To carry out basis registration, it is also necessary to know or estimate the $M-1$ remaining meaningful basis vectors in $\mathcal{P}$.
-In this case the three possibilities become: (a) the basis is known a priori; (b) the basis is estimated using expert knowledge, perhaps involving information beyond the landmarks themselves; or (c) hybrid Procrustes analysis.
-For details in Procrustes analysis, see, for example Dryden and Mardia (2016).
-By “hybrid”, we mean that (i) the normal direction to the midplane is estimated by combining an object with its reflection and using generalized Procrustes analysis (GPA), (ii) an “average shape” is found from all the objects using GPA, (iii) the midplane for the average shape is oriented so the basis directions have natural interpretations using expert knowledge, and (iv) each data configuration is aligned to the average shape using ordinary Procrustes analysis.
-The details are given in Appendix A.
-For the smile data application in Section 4, expert knowledge (scenario (b)) has been used to estimate the midplane with the stated interpretations for the coordinate axes as above in basis registration.
+Suppose there is an associated midplane $\mathcal{P}$ for every bilateral data object $X$.
+Ways in which $\mathcal{P}$ might be determined are discussed below.
+For data analysis, it is helpful to “register” $X$ using a rigid body motion so that after transformation $\bm{n}=\bm{e}_{1}$ and $b=0$, i.e., the midplane passes through the origin and the normal direction is given by the first coordinate axis, while still keeps the size-and-shape of $X$ unchanged.
+Figure 1: Coordinates system used for registration of human face used in our illustrative smile examples: with three principal planes.
+In the figure at top row, the $x$-$y$ plane is in green, the $y$-$z$ plane is in red and the $x$-$z$ plane is in blue.
+For the three sub-figures in the bottom row, the $x$-axis is in green, $y$-axis is in red and $z$-axis is in blue.
+In some settings it is possible to carry out a stronger version of registration than the axis registration, we will call basis registration, in which all the coordinate axes have a natural interpretation.
+In our illustrative smile examples, for 3D measurements of the human head, the midplane is known as the sagittal plane, and it is natural to require the coordinate axes to have the following interpretations:
+  • (a)
+    $x$-coordinate: left-right
+  • (b)
+    $y$-coordinate: down-up
+  • (c)
+    $z$-coordinate: back-front
+Figure 1 shows these coordinate axes as well as the three planes on a human face which have been used in pre-registration for our smile data.
+That is, the $x$-axis is positioned horizontally on the face (positive direction is left from the view point of the subject), the $y$-axis is placed vertically on the face (positive direction is upward), whereas the $z$-axis is positioned in the direction inwards and outwards of the face (positive direction is outward), see Figure 1.
+In general, to carry out basis registration, it is also necessary to know or estimate M1M-1 meaningful basis vectors in $\mathcal{P}$.
+In this case the three possibilities become: (a) the basis is known a priori; (b) the basis is estimated using expert knowledge; or (c) the basis is estimated using hybrid Procrustes analysis.
+For details of Procrustes analysis, see, for example, Goodall (1991) and Dryden and Mardia (2016).
+For our illustrative smile examples, expert knowledge (scenario (b)) has been used to estimate the midplane with the stated interpretations for the coordinate axes as above (Figure 1).
+It may be noted that there will be one more reflection before the registration in some data; specifically, for those cleft lip subjects in Smile Data 1 with the cleft on the right, we reflect in the midplane so that the cleft is consistently on the left.
 2.2 Elementary features
-Suppose the size-and-shape dataset has been basis registered about a midplane $\mathcal{P}$ that passes through the origin and is normal to the first coordinate axis.
+Suppose the size-and-shape dataset has been basis registered about a midplane $\mathcal{P}$ that passes through the origin and is normal to the first coordinate axis and the M1M-1 axes of $\mathcal{P}$ are all determined.
 Then it is possible to define various elementary features to describe the asymmetry of each data object.
 Paired and solo landmarks are treated separately.
-Let $K_{L}$ and $K_{R}$ denote the total number of landmarks on the left or right of $\mathcal{P}$.
-We have $K_{L}=K_{R}=K_{P}$.
-Let $k_{L}$ and $k_{R}$ denote the left and right indices for a typical landmark pair for a bilateral object $X$ and consider the $M$ coordinatewise signed elementary features for the landmark pairs,
+We begin with an idealized bilaterally symmetric object $X^{S}$.
+Each of the $K$ landmarks of $X^{S}$ will be either to the left of $\mathcal{P}$, to the right of $\mathcal{P}$, or on $\mathcal{P}$.
+The counts of landmarks in these subsets are respectively $K_{L}$, $K_{R}$, and $K_{S}$ where $K_{L}=K_{R}=K_{P}$ .
+For any $K\times M$ observation $X$ modeled on $X^{S}$, landmarks, identified by their indices, follow the prescription of $X^{S}$.
+Let $k_{L}$ and $k_{R}$ denote the left and right indices for a typical landmark pair from a bilateral object $X$ and consider the $M$ coordinatewise signed elementary features for the landmark pairs,
 $d[(k_{L},k_{R}),1]$ $=X[k_{L},1]+X[k_{R},1]$ (2.1)
 $d[(k_{L},k_{R}),m]$ $=X[k_{L},m]-X[k_{R},m],\quad m=2,\ldots,M.$ (2.2)
-That is, there is one feature for each coordinate and each landmark pair.
-So there are in total $MK_{P}$ coordinatewise features for all paired landmarks.
-For this specific $\mathcal{P}$, the first coordinates for the two landmarks in one pair have opposite signs, while other coordinates have the same sign.
-Further, for a symmetric object, the quantities in equations (2.1) and (2.2) should all be 0.
-Thus, taking the sum of the first coordinates and differences for other coordinates allow us to quantify the departure from symmetry.
+That is, there is one feature for each coordinate of each landmark pair, so there are in total $MK_{P}$ coordinatewise features.
+The first coordinates for the two landmarks in one pair have opposite signs, while other coordinates have the same sign with respect to $\mathcal{P}$.
+Note that this pattern is true for both axis and basis registrations.
+For $X^{S}$, the quantities in equations (2.1) and (2.2) should all be 0.
+Thus, the quantities given in (2.1) and (2.2) quantify the departure from symmetry.
 Similarly, if $k_{S}$ is a typical solo landmark, consider the single elementary feature
@@
 There are in total $K_{S}$ elementary features for solos.
-Under bilateral symmetry with $\mathcal{P}$ chosen as above, the first coordinate of a solo landmark will be 0, with no restriction on other coordinates.
+For $X^{S}$, the first coordinate of a solo landmark will be 0, with no restriction on other coordinates.
 Hence, a single feature as in (2.3) is adequate to describe the asymmetry of a solo landmark.
@@
 The elements of $\bm{d}$ will either be listed sequentially with a subscript index, i.e., $d_{j},\ j=1,\ldots,J_{\text{basis}}$, or using parentheses within square brackets, as in (2.1)-(2.3).
+Note that the vector $\bm{d}$ is a function of matrix $X$ and $\mathcal{P}$.
 When only axis registration is available, the coordinate information in (2.2) does not have a well-defined interpretation.
-This is because the coordinates except the first one of paired landmarks would be changed after rotation within $\mathcal{P}$, hence leads to changes in values of (2.2) even though the asymmetry information does not change.
-Note that the equation (2.1) will not change.
+This is because the coordinates except the first one of paired landmarks would be changed after rotation within $\mathcal{P}$, hence leads to changes in values of (2.2) even though the asymmetry information does not change (but note that equation (2.1) does not change).
 Further, at each landmark the coordinate information can be collected into a single number, called a landmark elementary feature which is defined per landmark,
-$d^{*}[(k_{L},k_{R})]=\left\{\sum_{m=1}^{M}d^{2}[(k_{L},k_{R}),m]\right\}^{1/2},$ (2.4)
-which does have a well-defined interpretation as the Euclidean distance between one landmark in a pair and the reflection of the other landmark in the pair.
-Rotation within $\mathcal{P}$ would not change this Euclidean distance.
-For a solo landmark, (2.3) is always well defined since the first coordinate of a solo remains unchanged after rotation within $\mathcal{P}$; hence it is both a coordinatewise and a landmark elementary feature.
-Thus, overall in the case of axis registration, a smaller set of $J_{\text{axis}}=K_{P}+K_{S}$ elementary features can be collected into a vector for further analysis.
-On the other hand, for $M=2$, there is no difference between axis and basis registration.
-Thus, equations (2.1)-(2.3) are all valid for both types of registrations in this case.
+$d^{*}[(k_{L},k_{R})]=\left\{\sum_{m=1}^{M}d^{2}[(k_{L},k_{R}),m]\right\}^{1/2},\ d^{*}[(k_{S})]=|X[k_{S},1]|,$ (2.4)
+where $d$’s are defined by (2.1) and (2.2).
 Another useful way to think about the elementary features is by using reflection.
-Let $X$ be a basis registered configuration, and let $X^{(\text{refl})}$ denote the reflection of $X$ about $\mathcal{P}$.
+Let $X^{(\text{refl})}$ denote the reflection of $X$ about $\mathcal{P}$.
 Computationally, this matrix is obtained by changing the sign of the first column of $X$ and interchanging the row indices for each landmark pair.
 Then the elementary features of $X$ can be obtained by comparing the elements of $X$ and $X^{(\text{refl})}$.
-If $X$ is bilaterally symmetric about $\mathcal{P}$, then $X=X^{(\text{refl})}$ and elementary features vanish, i.e.
-$\bm{d}=\bm{0}$.
-In many examples, such as the smile application studied later in this paper in Section 4, the direction of asymmetry, for example, left or right asymmetry, can vary between individuals and is not thought to be interested.
+If $X$ is bilaterally symmetric about $\mathcal{P}$, then $X=X^{(\text{refl})}$ and $\bm{d}=\bm{0}$.
+In many examples, such as the illustrative smile examples studied later in this paper in Section 5 and 6, the direction of asymmetry, for example, left or right asymmetry, can vary between individuals and is not thought interesting.
 We will focus on the absolute values of the elementary features.
@@
 where $J=J_{\text{basis}}$ or $J=J_{\text{axis}}$, as appropriate.
-In other words, only the extent of the asymmetry is interested.
+In other words, only the extent of the asymmetry is interesting.
 The elements of $\bm{a}$ can also be written in a similar way as elements of $\bm{d}$ in (2.1)-(2.4).
 In the rest of the paper, each configuration $X$ is reduced to an absolute elementary feature vector $\bm{a}$ for further numerical and statistical analysis.
-2.3 A Simple Illustrative Example
-To illustrate the construction of the vector $\bm{a}$, consider two configurations in $M=2$ dimensions with $K=4$ landmarks given by
-$X_{1}=\begin{pmatrix}-1&\phantom{-}0\\ \phantom{-}0&\phantom{-}1\\ \phantom{-}1&\phantom{-}0\\ \phantom{-}0&-1\end{pmatrix},\quad X_{2}=\begin{pmatrix}-0.95&\phantom{-}0.36\\ -0.28&\phantom{-}2.11\\ \phantom{-}0.99&\phantom{-}0.54\\ -0.31&-1.37\end{pmatrix}.$ (2.6)
-Suppose the landmarks consist of one landmark pair, $(k_{L},k_{R})=(1,3)$ and two solos, $k_{S}=2,4$.
-Hence $K_{P}=1$ and $K_{S}=2$.
-Further, suppose both configurations are treated as basis registered with $\mathcal{P}$ normal to the first coordinate axis, i.e.
-$\mathcal{P}$ is the $y$-axis.
+Note that $\bm{a}$ is also a function of matrix $X$ and $\mathcal{P}$.
+2.3 Example 1
+We give here a simple example to illustrate our basic notation and construction of the vector $\bm{a}$ and related quantities.
+Consider two configurations in $M=2$ dimensions with $K=4$ landmarks given by
 […]
@@
 The left and right figures in the bottom row show the original objects (in black solid lines) together with their respectively reflections $X_{1}^{\text{(refl)}}$ and $X_{2}^{\text{(refl)}}$ (in red dotted lines) respectively, in order to illustrate the coordinatewise elementary feature vector.
+Therefore, there is one landmark pair $(k_{L},k_{R})=(1,3)$ and two solos, $k_{S}=2,4$, so $K_{P}=1$ and $K_{S}=2$.
+Further, suppose both configurations are treated as basis registered with $\mathcal{P}$ normal to the first coordinate axis, i.e.
+$\mathcal{P}$ is the $y$-axis.
+The configurations $X_{1}$ and $X_{2}$ are shown in the left and right figures in the top row of Figure 2 respectively, whereas the bottom row of Figure 2 show their corresponding reflections, where the left is for $X_{1}$ and $X_{1}^{\text{(refl)}}$ and right is for $X_{2}$ and $X_{2}^{\text{(refl)}}$.
+Note that $X_{1}$ is bilaterally symmetric, with $\bm{a}=\bm{0}$.
+Table 1 gives the $J_{\text{basis}}=4$ elements of the vector $\bm{a}$ for $X_{2}$.
+The two elements for the landmark pair are listed first, and the two elements for the solos listed last, with
+$\bm{a}=(0.04,\ 0.18,\ 0.28,\ 0.31)^{T}$
+for $X_{2}$ .
+It can be seen visually in Figure 2 that for $X_{2}$, landmarks 2 and 4 deviate more from the $y$-axis than the landmark pair (1, 3) differs from symmetry.
+Hence the last two components of $\bm{a}$ are larger than the first two components.
 Table 1: Example in Section 2.3. Elements of the absolute elementary feature vector $\bm{a}\in\mathbb{R}^{4}$ for $X_{2}$ with $K=4$ landmarks in $M=2$ dimensions.
@@
 3 Hypothesis tests
-Consider a dataset of $N$ basis registered configurations observed on two groups of individuals, where the first $N_{1}$ configurations belong to one group and the last $N_{2}$ configurations belong to the other group, $N=N_{1}+N_{2}$, where it is thought that the first group of configurations might be less asymmetric than the second.
+Consider a dataset of $N$ basis registered configurations observed on two groups of individuals, where the first $N_{1}$ configurations belong to one group (Group 1) and the last $N_{2}$ configurations belong to the other group (Group 2), $N=N_{1}+N_{2}$, where it is thought that the first group of configurations might be more symmetric than the second.
 One way to think about this question is to construct a one-sided hypothesis test of
-$\begin{split}&H_{0}:\text{ the two groups have the same distribution of asymmetry vs.}\\ &H_{1}:\text{ Group 1 is less asymmetric than Group 2.}\end{split}$ (3.1)
-We will give two general approaches to construct a test statistic: (a) combine-then-compare, and (b) compare-then-combine. These two approaches will be explored in the next subsections.
-To specify the null and alternative hypotheses explicitly, some notations are needed. Suppose the data take the form of $N$ registered configurations
+[…]
+For the paired data, we change the Group 1 and Group 2 in the above hypotheses to experimental and control groups respectively, where same subjects are contained in the two paired groups.
+We will give two general approaches to construct a test statistic in order to compare the symmetries between the two groups: (a) combine-then-compare, where univariate test methods are used, and (b) compare-then-combine, where we form the hypotheses under union intersection test framework.
+These two approaches will be explored in the next subsections.
+Suppose the data take the form of $N$ registered configurations
 $X_{n},\ n=1,\ldots,N.$ (3.2)
@@
 Then the hypotheses given in equation (3.1) can be reformulated as:
-$\begin{split}&H_{0}:\text{ $\bm{a}_{n}\sim F$ i.i.d.
-for all $n$ vs.}\\ &H_{1}:\text{ $\bm{a}_{n}\sim F_{1}$ i.i.d.
-for $n=1,\dots,N_{1}$, $\bm{a}_{n}\sim F_{2}$ i.i.d.
-for $n=N_{1}+1,\dots,N$,}\end{split}$ (3.3)
-where $F$, $F_{1}$, $F_{2}$ are multivariate cumulative density functions (c.d.f.s) and $F_{1}$ is stochastically smaller than $F_{2}$ (that is, $F_{1}(\bm{a})\geq F_{2}(\bm{a})$ for all $\bm{a}$ with at least one strict inequality), see, for example Stoyan (1983).
+[…]
+where $F$, $F_{1}$, $F_{2}$ are multivariate cumulative density functions (c.d.f.s) and $F_{1}$ is stochastically smaller than $F_{2}$ (that is, $F_{1}(\bm{a})\geq F_{2}(\bm{a})$ for all $\bm{a}$ with at least one strict inequality, see, for example Stoyan (1983)).
 In other words, the distribution of $F_{2}$ is shifted towards right comparing with $F_{1}$, which indicates more asymmetries.
-We concentrate in this section on t-test and Mann-Whitney U test. In Section 4, we further discuss some alternative univariate tests as well as multivariate tests in relation to the smile data.
+Under the paired data setting, $\bm{a}_{n}$ for $n=1,\dots,N_{1}$ represent the experimental data and $\bm{a}_{n}$ for $n=N_{1}+1,\dots,N$ represent the control data, where $N=2N_{1}$.
+We concentrate in this section on t-test and Mann-Whitney U test.
+In Section 5, we further discuss some alternative univariate tests as well as multivariate tests in relation to the application of Smile Data 1.
 3.1 Testing strategy: combine-then-compare
@@
 $u=\phi(\bm{a}),$ (3.4)
-say. Here the function $\phi(\bm{a})$ is assumed to be monotonically increasing in the sense that it is monotonically increasing in each component $a_{j}$ when the other elements are held fixed.
+say. Here the function $\phi(\bm{a})$ is assumed to be monotonically increasing in each component $a_{j}$ when the other elements are held fixed.
 Let $\psi(a)$ be a monotonically nonegative increasing function of a scalar argument $a\geq 0$. Then one way to define $\phi(\bm{a})$ is as an additive composite asymmetry score
@@
   • (b)
-    greater weights near the midplane,
+    greater weights near the midplane or away from the midplane, or
   • (c)
-    greater weights away from the midline, or
-  • (d)
     different weights for pairs and solos.
-For example, in cases (b) and (c) a weight might depend on $|X[k_{R},1]-X[k_{L},1]|$ for a landmark pair, with solos treated separately.
-The choices of weights in general depend on prior belief of the extent of asymmetry information carried by each landmark, for example, if it is known a priori that landmarks far away from $\mathcal{P}$ can carry important asymmetry information than landmarks near $\mathcal{P}$, then more weights should be assigned to the farther landmarks.
-For smile example, this information is available so greater weights will be given near the midplane, see Section 4.
-The choice of function $\psi$ governs the relative influence of different magnitudes of an elementary feature.
+For example, in case (b), a weight might depend on $|X[k_{R},1]-X[k_{L},1]|$ for a landmark pair, with solos treated separately.
+The choices of weights in general depend on prior belief of the extent of asymmetry information carried by each landmark, for example, if it is known a priori that landmarks near $\mathcal{P}$ can carry more important asymmetry information than landmarks far away from $\mathcal{P}$, then greater weights should be assigned to the nearer landmarks.
+For Smile Data 1, this information is available so greater weights will be given near the midplane, see Section 5.
+Another way to give some elementary features greater prominence than others is through the choice of the functions $\psi$.
 For example, the quadratic function $\psi(a)=a^{2}$ is more sensitive to outlying features than the linear function $\psi(a)=a$.
-We write below explicitly for two such scores with equal weights and give them names $L_{1}$ and $L_{2}$ statistics respectively.
-[…]
-[…]
-In the above equations, we use the two expressions for elements in $\bm{d}$ and $\bm{a}$ introduced in Section 2.2.
-Bock and Bowman (2006) have proposed an asymmetry score with weights $w_{j}=1$ for landmark pairs and weights $w_{j}=2$ for solos and $\psi(a_{j})=a_{j}^{2}$ in equation (3.5), namely,
-$\sum_{(k_{L},k_{R})}\sum_{m=1}^{M}d[(k_{L},k_{R}),m]^{2}+2\sum_{k_{S}}d[(k_{S})]^{2}.$ (3.8)
-For our work, we have selected an equal weights for $\phi_{L_{1}}$ to have some similarity to the asymmetry score of Patel et al. (2023); their score is given by
-$\phi_{L_{1}}^{*}(\bm{a})=\sum_{(k_{L},k_{R})}d^{*}[(k_{L},k_{R})]+\sum_{k_{S}}|d[(k_{S})]|,$ (3.9)
-with weights equal to 1 for both pairs and solos. The $d^{*}[(k_{L},k_{R})]$ is defined in equation (2.4). Also, we can define a similar score as follows using $L_{2}$ distances
-$\phi_{L_{2}}^{*}(\bm{a})=\sum_{(k_{L},k_{R})}d^{*}[(k_{L},k_{R})]^{2}+\sum_{k_{S}}|d[(k_{S})]|^{2}.$ (3.10)
-It can be shown using (2.4) that this equation (3.10) reduces to (3.7). That is, we have
-$\phi^{*}_{L_{2}}(\bm{a})=\phi_{L_{2}}(\bm{a}).$
-Note that, we can generalize $\phi^{*}(\bm{a})$ in the same way as $\phi(\bm{a})$ with weights.
-Example in Section 2.3 continued.
+We write below explicitly two such scores with equal weights and give them names $L_{1}$ and $L_{2}$ statistics respectively.
+$\text{$L_{1}$ statistic: }\phi_{L_{1}}(\bm{a})=\sum_{j=1}^{J}a_{j}=\sum_{(k_{L},k_{R})}\sum_{m=1}^{M}|d[(k_{L},k_{R}),m]|+\sum_{k_{S}}|d[(k_{S})]|,$ (3.6)
+$\text{$L_{2}$ statistic: }\phi_{L_{2}}(\bm{a})=\sum_{j=1}^{J}a_{j}^{2}=\sum_{(k_{L},k_{R})}\sum_{m=1}^{M}d[(k_{L},k_{R}),m]^{2}+\sum_{k_{S}}d[(k_{S})]^{2},$ (3.7)
+where the $d$’s are given in (2.1)-(2.3) and $\bm{a}$, $a_{j}$ are given in (2.5).
+We note that Bock and Bowman (2006) have proposed an asymmetry score which is proportional to $\phi_{L_{2}}$, that is
+$\frac{2}{K}\phi_{L_{2}}(\bm{a}).$ (3.8)
+Further, Patel et al. (2023) uses the score given by
+$\phi_{L_{1}}^{*}(\bm{a})=\sum_{(k_{L},k_{R})}d^{*}[(k_{L},k_{R})]+\sum_{k_{S}}d^{*}[(k_{S})],$ (3.9)
+The $d^{*}[(k_{L},k_{R})]$ and $d^{*}[(k_{S})]$ are defined in equation (2.4). Also, by squaring each element in the two sums in (3.9) and using (2.4), we recover $\phi_{L_{2}}$:
+$\phi_{L_{2}}^{*}(\bm{a})=\sum_{(k_{L},k_{R})}d^{*}[(k_{L},k_{R})]^{2}+\sum_{k_{S}}|d^{*}[(k_{S})]|^{2}=\phi_{L_{2}}(\bm{a}).$ (3.10)
+Note that we can generalize $\phi^{*}(\bm{a})$ in the same way as $\phi(\bm{a})$ with weights.
+Instead of considering asymmetries with respect to each landmark as in (3.9), (3.6) and (3.7) view all features in $\bm{a}$ as a whole and take the $L_{1}$ and $L_{2}$ norms.
+For the subjects which are more symmetric, their corresponding $\bm{a}_{n}$ should be close to the origin.
+Hence, the $L_{1}$ and $L_{2}$ distances between $\bm{a}_{n}$ with $\bm{0}\in\mathbb{R}^{J}$ are expected to be smaller for control subjects.
+Example 1 continued.
 Recall the asymmetric square $X_{2}$ defined in equation (2.6).
@@
 Note that the values of
-$d^{*}[(1,3)]=\sqrt{(X_{2}[1,1]+X_{2}[3,1])^{2}+(X_{2}[1,2]-X_{2}[3,2])^{2}}=0.18$
+$\begin{split}d^{*}[(1,3)]=\sqrt{(X_{2}[1,1]+X_{2}[3,1])^{2}+(X_{2}[1,2]-X_{2}[3,2])^{2}}=0.18\end{split}$
 and $a[(2)]=0.28$, $a[(4)]=0.31$ (Table 1) are less than 1, so taking squares of them in $\phi_{L_{2}}$ leads to smaller values.
 Note that here we have scaled $\phi^{*}_{L_{1}}$ by a factor of half, so unscaled $\phi^{*}_{L_{1}}=0.78$ which is similar to the value of $\phi_{L_{1}}=0.81$.
+Table 2: The two composite asymmetry scores defined in equation (3.5) and equation (3.9) are computed on the configuration $X_{2}$.
 $d^{*}[(1,3)]$ (2.4) $\phi_{L_{1}}(\bm{a})$ (3.6) $\phi^{*}_{L_{1}}(\bm{a})$ (3.9) $\phi_{L_{2}}(\bm{a})$ (3.7)
 $X_{2}$ 0.18 0.81 0.39 0.21
-Table 2: The two composite asymmetry scores defined in equation (3.5) and equation (3.9) are computed on the configuration $X_{2}$.
 Once a composite asymmetry score $\phi(\bm{a})$ has been chosen, let
@@
 denote the composite asymmetry scores for the data in (3.2).
-Let
+Suppose $u_{n}$ for $n=1,\ldots,N_{1}$ and $n=N_{1}+1,\ldots,N$ are realizations from random variables $U_{1}$ and $U_{2}$ respectively, and let $g=1,2$ denote the control and cleft groups. Let
 $\mu_{g}=E\{U_{g}\},\ \sigma^{2}_{g}=\text{var}\{U_{g}\}$ (3.12)
-denote the expectation and variance for a random variable $U_{g}\sim F_{g}$, where $g=1,2$ labels one of the two groups. A simpler version of hypotheses in (3.1) is
+denote the expectation and variance for a random variable $U_{g}\sim F_{g}$ (should not be confused with the distribution of the vector $\bm{a}$ in Section 3).
+A simpler version of hypotheses in (3.1) is
 $H_{0}:\mu_{1}=\mu_{2}\text{ vs. }H_{1}:\mu_{1}<\mu_{2}$ (3.13)
@@
   • (a)
-    the standard two-sample t-test, which assumes a common variance for the two groups ($\sigma_{1}^{2}=\sigma_{2}^{2}$);
+    Parametric test: the standard two-sample t-test, which assumes $\sigma_{1}^{2}=\sigma_{2}^{2}$;
   • (b)
-    the Mann-Whitney U test, also known as the Wilcoxon rank sum test (Mann and Whitney, 1947).
-The Mann-Whitney U test is nonparametric so that the test results are invariant under monotonic transformations of the $u_{n}$.
+    Non-parametric test: the Mann-Whitney U test, also known as the Wilcoxon rank sum test (Mann and Whitney, 1947).
 In case (a), the significance of the test statistic under $H_{0}$ can be computed either analytically (assuming normality) or using a bootstrap approximation.
@@
 Mann-Whitney U test is considered where the normality assumption on composite asymmetry scores $u_{n}=\phi(\bm{a}_{n})$ may fail to hold.
-An advantage of the combine-then-compare approach is that it enables us to accumulate evidence from different features.
-If there is a small amount of asymmetry on many different elementary features, then the composite score will have a large value.
+Paired data The test strategies described above can be applied to the paired data by making necessary modification, namely we need to use the paired test in place of the two-sample test for both the parametric and non-parametric cases.
+Specifically, paired t-test and Wilcoxon signed-rank test are used for the parametric and non-parametric cases respectively.
+An advantage of the combine-then-compare approach is that it enables us to accumulate evidence from different features, so if there is a small amount of asymmetry on many different elementary features, then the composite score will have a large value.
 Further, this approach enables the two groups to be compared visually with stem and leaf plots and in particular, any overlap between the two groups can be easily assessed.
-Applications of this approach are in Section 4
+Applications of this approach are in Section 5.2.1, 6.2.1 and 6.3.1.
 3.2 Testing strategy: compare-then-combine
@@
 The most extreme of these separate test statistics is then used as an overall test statistic.
-Before we give the details, the hypotheses in equation (3.1) need to be formulated in the union intersection test style (as explained in Section 1) as following:
+Before we give the details, the hypotheses in equation (3.1) need to be formulated in the Union Intersection Test (UIT) style:
 $H_{0}=\bigcap_{j=1}^{J}H_{0}^{j}\text{ vs }H_{1}=\bigcup_{j=1}^{J}H_{1}^{j},$ (3.14)
 where $H_{0}^{j}$: the mean of $j$th unsigned elementary feature between the two groups is the same, whereas $H_{1}^{j}$ is that the mean of group 2 is larger than mean of group 1, for $j=1,\dots,J$.
-For each choice $j$ of an unsigned elementary feature, construct a statistic to compare the two groups, for example, a two-sample t-statistic or a Mann-Whitney U statistic.
+For each choice $j$ of an unsigned elementary feature, we construct a statistic to compare the two groups, for example, a two-sample t-statistic or a Mann-Whitney U statistic.
 Denote the resulting statistics by $v_{j},\ j=1,\ldots,J$.
@@
 $v_{j}=(\bar{a}_{j}^{(1)}-\bar{a}_{j}^{(2)})/s_{j}.$ (3.15)
-where $\bar{a}_{j}^{(g)},\ g=1,2$ are the sample means of the $j$th unsigned elementary feature in the two groups and $s^{2}_{j}$ is the pooled within-group variance.
-An overall UIT statistic to test $H_{0}$ vs $H_{1}$ in equation (3.14) can be defined by taking the maximum of these separate statistics (Boyett and Shuster, 1977),
-$V=\max_{j=1,\ldots,J}v_{j}.$ (3.16)
-The critical points of $V$ under the null hypothesis can be computed by the bootstrap.
-Note that the landmarks are correlated with each other, so the $d_{j}$ are not independent for $j=1,\ldots,J$ here, so we cannot derive the distribution of $V$ theoretically.
-Hence, bootstrap is used for estimating the critical points of $V$.
-If there exists $v_{j}$ which exceeds the critical values of $V$, then we not only know that $H_{0}$ is rejected, but also know that the landmark corresponding to such $v_{j}$ is also important.
-Our main aim is somewhat different from Boyett and Shuster (1977): we aim at selecting important landmarks while Boyett and Shuster (1977), attempt to estimate a $p$-value.
-The bootstrap procedure used by us is also different from their who have used:
-  1. 1.
-    A set $E$ containing all permutations of the dataset is constructed.
-  2. 2.
-    Sample with replacement from this set $E$, i.e. sample permutations of the data. The permuted samples constitute the bootstrap resampled dataset.
-  3. 3.
-    Test statistic $V$ is computed on the resampled dataset.
-On the other hand, we directly sample from the whole dataset as following:
-  1. 1.
-    Sample $N$ samples with replacement from $\{\bm{a}_{n}\}_{n=1}^{N}$.
-  2. 2.
-    Compute $v_{j}$ and $V$ on resampled dataset.
-An advantage of this compare-then-combine approach versus combine-and-compare is that if the null hypothesis is rejected, then simultaneous confidence intervals can be constructed to identify those features on which the two groups significantly differ.
-The selection procedure in spirit is similar to Tukey’s method (Tukey, 1949) in ANOVA though in ANOVA a separate test statistic is used for each pairwise difference between main effects.
-Here, a separate test statistic for each $j$ is more appropriate since all features in unsigned elementary feature vector $\bm{a}$ have different biology.
-An illustration is given below for the smile data.
-4 Analyses of the Smile Data
-We now give smile data taken from Patel et al.
-(2023) which have been pre-registered.
-The data is collected using the Di4D system.
-When a subject’s face is placed in front of the camera, the $x$-axis is positioned horizontally on the face (positive direction is left), the $y$-axis is placed vertically on the face (positive direction is upward), whereas the $z$-axis is positioned in the direction inwards and outwards of the face (positive direction is outward), see Figure 1.
-The coordinates given in Section 2.1 can be related to these coordinate axes as follows:
-coordinate 1: left-right $x$
-coordinate 2: down-up $y$
-coordinate 3: back-front $z$.
-Our smile data contains $N=25$ subjects, of which $N_{1}=12$ are control subjects and $N_{2}=13$ are cleft subjects.
+where $\bar{a}_{j}^{(g)},\ g=1,2$ are the sample means of the $j$th unsigned elementary feature in the two groups and $s^{2}_{j}$ is the corresponding pooled within-group variance.
+Following Boyett and Shuster (1977), an overall UIT statistic to test $H_{0}$ vs $H_{1}$ in equation (3.14) can be defined by taking the maximum of these separate statistics,
+$V=\max_{j=1,\ldots,J}v_{j},$ (3.16)
+and Boyett and Shuster (1977) compute the critical points of $V$ under the null hypothesis using bootstrap.
+The landmarks are correlated with each other, so the $d_{j}$ are not independent for $j=1,\ldots,J$ here, so we cannot derive the distribution of $V$ theoretically.
+Hence, the bootstrap has been used for estimating the critical points of $V$.
+If there exists $v_{j}$ which exceeds the critical value of $V$ at a pre-determined significance level, then we not only know that the corresponding $H_{0}^{j}$ is rejected (hence the overall $H_{0}$ in (3.14) is also rejected), but also know that the landmark corresponding to such $v_{j}$ is important.
+Our main aim is somewhat different from Boyett and Shuster (1977): we aim at discovering important landmarks while Boyett and Shuster (1977) focuses on estimating a $p$-value.
+Also, the two bootstrap procedures are slightly different.
+Their procedure is as follows:
+  1. (a)
+    Construct a set $E_{n}$ containing all random samples of size $n$ taken without replacement from origin dataset $\{\bm{a}_{i}\}_{i=1}^{N}$.
+  2. (b)
+    Sample with replacement $B$ times from this set $E_{n}$ to form the bootstrap resampled dataset.
+  3. (c)
+    Compute the test statistic $\{V^{(1)},\dots,V^{(B)}\}$ on the resampled dataset.
+  4. (d)
+    Compare the test statistic $V$ computed on first $n$ observed data $\{\bm{a}_{i}\}_{i=1}^{n}$ to $\{V^{(1)},\dots,V^{(B)}\}$.
+The procedure given in Boyett and Shuster (1977) does not perform bootstrap directly on the data, and only part of the data is used in either $E_{n}$ or the observed statistic, which is probably to reduce the computational budget.
+In contrast, since our data size is not large and the computational budget is far more than enough, we directly sample with replacement from the whole dataset with $n=N$ and use all data to compute the observed statistic.
+Using all data is very likely to improve the final result.
+Our procedure is as follows:
+  1. (a)
+    Sample $N$ samples with replacement from $\{\bm{a}_{i}\}_{i=1}^{N}$ and repeat this step $B$ times.
+  2. (b)
+    Compute $\{V^{(1)},\dots,V^{(B)}\}$ on resampled dataset.
+  3. (c)
+    Compare the test statistic $V$ computed on all observed data $\{\bm{a}_{i}\}_{i=1}^{N}$ to $\{V^{(1)},\dots,V^{(B)}\}$.
+For compare-then-combine approach, we use a similar selection procedure to Tukey’s method (Tukey, 1949) in ANOVA though in ANOVA a separate test statistic is used for each pairwise difference between main effects.
+Here, a separate test statistic for each $j$ is more appropriate since the different features in the unsigned elementary feature vector $\bm{a}$ measure different features of the face.
+An illustration is given below in Section 5.2.2 and 6.3.2.
+Paired data The approach given in this section can be extended to the paired data scenario with the following changes applied to the bootstrap procedure.
+  1. (a)
+    Compute $a^{(2)}_{nj}-a^{(1)}_{nj}$ for each subject $n$ and each feature $j$, where $a^{(1)}_{nj}$ and $a^{(2)}_{nj}$ are the feature $j$ of vectors $\bm{a}_{n}^{(1)}$ and $\bm{a}_{n}^{(2)}$ for subject $n$ from Group 1 and 2 respectively.
+  2. (b)
+    Assign +1+1 or 1-1 to each $\bm{a}^{(2)}_{n}-\bm{a}^{(1)}_{n}$ randomly with equal probability, which forms the permuted sample. Repeat this step $B$ times.
+  3. (c)
+    Compute $t$-values on the permuted sample and record the largest $t$-value.
+An illustration of this compare-then-combine approach is given below for the Smile Data 2 in Section 6.2.2.
+4 Meta-Analysis Methods
+We can consider the meta-analysis, which carries out simultaneous inference and also belongs to the compare-then-combine approach category.
+In particular in meta-analysis, we can obtain an overall $p$-value by combining the $p$-values from separate tests.
+We now give an overview for meta-analysis.
+Suppose we have $N$ basis registered configurations, where the first $N_{1}$ configurations are from Group 1 and the rest $N_{2}$ configurations are from Group 2.
+Absolute elementary feature vectors $\bm{a}_{1},\dots,\bm{a}_{N}\in\mathbb{R}^{J}$ defined in equation (2.5) can be constructed on each configuration.
+Let $\bm{a}^{(g)}$ be the random vector associated to the population of group $g$ and let $a_{j}^{(g)}$ denote feature $j$ of vector $\bm{a}^{(g)}$, where $g=1,2$ for Group 1 and Group 2 respectively and $j=1,\dots,J$.
+Consider the $J$ individual tests with corresponding hypotheses
+$\begin{split}H_{0}^{j}:a_{j}^{(1)}=a_{j}^{(2)},\ \ H_{1,L}^{j}:a_{j}^{(1)}<a_{j}^{(2)},\ \ H_{1,R}^{j}:a_{j}^{(1)}>a_{j}^{(2)},\ \ H_{1}^{j}:a_{j}^{(1)}\neq a_{j}^{(2)},\end{split}$
+for $j=1,\dots,J$.
+The $p$-values $\tilde{p}_{j}$ and $1-\tilde{p}_{j}$ correspond to the left and right alternative hypotheses $H_{1,L}^{j}$ and $H_{1,R}^{j}$ respectively.
+The two-sided $p$-value $p_{j}=2\min(\tilde{p}_{1},1-\tilde{p}_{j})$ corresponds to the two-sided alternative hypothesis $H_{1}^{j}$.
+The way of computing $\tilde{p}_{j}$ and $p_{j}$ depends on the choice of the univariate test.
+One of such methodology to combine individual $p$-values is the well-known Fisher’s method (Fisher, 1932). The Fisher’s combination function is given as:
+$P_{L}=-2\sum_{j=1}^{J}\log(\tilde{p}_{j}),\ P_{R}=-2\sum_{j=1}^{J}\log(1-\tilde{p}_{j}),P=-2\sum_{j=1}^{J}\log(p_{j}),$ (4.1)
+which correspond to the left, right and two-sided overall alternative hypotheses.
+$P_{L}$, $P_{R}$ and $P$ all follow $\chi^{2}_{2J}$-distribution.
+Hence, a final $p$-value is obtained by comparing them with the tail of the $\chi^{2}_{2J}$-distribution.
+For the Fisher’s method, $P_{L}$ is used in our illustrative smile examples but not $P$ or $P_{R}$, since we would like to test whether the values of absolute elementary features from Group 1 are smaller than the feature values from Group 2.
+In other words, we are interested in the left-sided tests.
+However, the right-sided and two-sided tests could be relevant for some other practical examples.
+There is another combination function based on Pearson (1934) which is given as
+$Q=\max(P_{L},P_{R}),$ (4.2)
+which does not follow a $\chi^{2}$-distribution.
+Hence, randomization tests and permutation tests are used to estimate the overall $p$-value in two-sample and paired data cases respectively.
+For the randomization tests, the procedures are given as below:
+  1. (a)
+    Sample without replacement from the pooled dataset of the two groups for $B$ times.
+  2. (b)
+    Perform individual two-sample tests on the resampled data and obtain $p$-values $\tilde{p}_{j}^{(b)}$ and $1-\tilde{p}_{j}^{(b)}$, for $j=1,\dots,J$ and $b=1,\dots,B$.
+  3. (c)
+    Compute the overall $p$-values defined in equation (4.1) and hence $Q$ defined in (4.2).
+The procedures for permutation tests are given as the following:
+  1. (a)
+    Shuffle the subjects randomly, such as change the subjects between experimental group to control group while keeping the paired structure. Repeat this step for $B$ times.
+  2. (b)
+    Perform individual paired tests to obtain $p$-values $\tilde{p}_{j}^{(b)}$ and $1-\tilde{p}_{j}^{(b)}$, for $j=1,\dots,J$ and $b=1,\dots,B$.
+  3. (c)
+    Compute the overall $p$-value defined in equation (4.1) and hence $Q$ defined in (4.2).
+We use $B=10000$ in both randomization and permutation tests.
+The choice between Fisher’s method and Pearson’s method depend on the problem at hand and Owen (2009) provides a review of both methods.
+In particular, if we suspect there are large deviations from individual null hypotheses, then Pearson’s method has higher power than Fisher’s method; the paper also shows that the Pearson’s method is admissible.
+The difference between the feature selection (see Section 3.2) and the Fisher’s and Pearson’s methods is that in the first approach, we record the test statistics resulted in each single test and use an overall test statistics to combine them together.
+Hence, obtaining a final $p$-value or carrying out feature selection are both possible.
+In the second approach, we combine the $p$-values obtained from each individual test together via a pre-defined combination function and calculate a final $p$-value.
+We cannot proceed the feature selection by using this method.
+The applications of the Fisher’s and Pearson’s methods on the two illustrative smile examples are given in Section 5.2.2, 6.2.2 and 6.3.2 respectively.
+5 Analyses of the Cleft Lip Smile Data
+5.1 Details
+We now illustrate the methods using the cleft lip smile data.
+The data is collected using the Di4D system (Patel et al., 2023) as described in Section 2 and it has been pre-registered using the basis registration.
+This data contains $N=25$ subjects, of which $N_{1}=12$ are control subjects and $N_{2}=13$ are cleft subjects.
 There are $K=24$ landmarks on the lip periphery.
@@
 The $M=3$ dimensional coordinates of these landmarks have been extracted at three frames: first (closed lip), middle (middle of the smile) and last frame (maximum open lip smile).
+We have $J=MK_{P}+K_{S}=35$.
 Let $g=1,2$ denote the control and cleft groups, respectively.
-Figure 3: Landmark indices for $x$-$y$ coordinates on the lip periphery of a control subject at first frame.
-Table 3: Indices of landmark pairs and solos in Figure 3 on the lip periphery.
+Figure 3: Landmark indices for $x$-$y$ coordinates on the lip periphery of a control subject at first frame from Smile Data 1.
+Table 3: Indices of landmark pairs and solos in Figure 3 on the lip periphery for the Smile Data 1.
 Landmark notation Indices in Figure 3
@@
 $k_{S}$ 7, 19
-In the rest of this section several test statistics will be presented in two parts; combine-then-compare approach and compare-then-combine approach.
-Combine-then-compare approach using (3.9).
+In the rest of this section several test statistics as described in Section 3 will be presented under two headings “combine-then-compare approach” and “compare-then-combine approach”.
+5.2 Test Results for Smile Data 1
+5.2.1 Application of the Combine-then-compare Approach
 Patel et al.
@@
 Further, the one-sided t-test indicates that there is more asymmetry for cleft lip subjects versus controls.
-Note that here the $p$-value is the smallest at the first frame, which suggests that the two groups are the most different at the start.
-Table 4: Mean, variance, t-values and $p$-values for $\phi^{*}_{L_{1}}(\bm{a})$ (equation (3.9)) for cleft and control subjects at first, middle and last frames (Patel et al. (2023)).
-Cleft Control One-sided test
-mean sd mean sd $t$-values $p$-values
-First 1.78 0.46 1.12 0.41 3.86 0.0004(***)
-Middle 1.96 0.70 1.46 0.47 2.14 0.02(*)
-Last 2.14 0.80 1.55 0.43 2.34 0.01(**)
+Note that here the $p$-value is the smallest at the first frame, which suggests that the two groups are the most different at the beginning of the smile.
+Table 4: Mean, variance, t-values and $p$-values for $\phi^{*}_{L_{1}}(\bm{a})$ using Smile Data 1 (equation (3.9)) for cleft and control subjects at first, middle and last frames (Patel et al.
+(2023)).
+Cleft Control One-sided test Two-sided test
+mean sd mean sd $t$-values $p$-values $p$-values
+First 21.35 5.54 13.39 5.11 -3.72 0.0006(***) 0.0011(**)
+Middle 23.57 8.37 17.35 5.87 -2.13 0.02(*) 0.04(*)
+Last 25.65 9.66 18.22 5.21 -2.36 0.01(*) 0.03(*)
 (*) = significant at the $5\%$ significance level. (**) = significant at the $1\%$ significance level. (***) = significant at the $0.1\%$ significance level.
-Combine-then-compare approach using (3.6) and (3.7).
-We use the scores $\phi_{L_{1}}(\bm{a})$ and $\phi_{L_{2}}(\bm{a})$ defined in equations (3.6) and (3.7) respectively.
-Instead of considering asymmetries with respect to each landmark as in (3.9), we view all features in $\bm{a}$ as a whole and take the $L_{1}$ and $L_{2}$ norms.
-For the subjects which are more symmetric, their corresponding $\bm{a}_{n}$ should be close to the origin.
-Hence, the $L_{1}$ and $L_{2}$ distances between $\bm{a}_{n}$ with $\bm{0}\in\mathbb{R}^{J}$ should be smaller for control subjects.
+Then, we use the scores $\phi_{L_{1}}(\bm{a})$ and $\phi_{L_{2}}(\bm{a})$ defined in equations (3.6) and (3.7) respectively.
 Several one-sided two-sample Mann-Whitney U tests have been performed and the results are shown in Table 5.
 For the last two rows in the table, the weighted $\phi_{L_{1}}(\bm{a})$ and weighted $\phi_{L_{2}}(\bm{a})$ are used and $w_{j}$ are chosen in an adaptive way as the following:
-  1. 1.
-    Compute the sample mean shape on the augmented data, which contains basis registered original configurations and their reflections through midplane.
+  1. (a)
+    Compute the sample mean shape on the doubled dataset, which contains basis registered original configurations and their reflections through midplane.
 Since the data has been pre-registered, the sample mean shape is simply the arithmetic mean for each landmark.
-  2. 2.
+  2. (b)
     The Euclidean distance within each landmark pair for this mean shape is computed and its reciprocal is used as the weight.
-The weights are the same among these subjects from both groups.
-The weights for solo landmarks are the unit length based on the scale of data.
+The weights are the same for all 25 subjects.
+The weights for solo landmarks are the unit length based on the scale of data ($w_{j}=1$ in (3.6) and (3.7)).
 Different landmarks carry different amount of asymmetry information.
@@
 Thus, we give higher weights to landmarks near $\mathcal{P}$ and smaller weights on landmarks faraway from the central.
-Table 5: $p$-values from one-sided two-sample Mann-Whitney U tests using $\phi_{L_{1}}(\bm{a})$ (3.6) and $\phi_{L_{2}}(\bm{a})$ (3.7).
+Table 5: $p$-values from one-sided two-sample Mann-Whitney U tests using $\phi_{L_{1}}(\bm{a})$ (3.6) and $\phi_{L_{2}}(\bm{a})$ (3.7) based on Smile Data 1.
 Composite asymmetry score First frame Middle frame Last frame
-$\phi_{L_{1}}(\bm{a})$ $9.69\times 10^{-5}$(***) 0.026(*) 0.009(**)
-$\phi_{L_{2}}(\bm{a})$ 0.0002(***) 0.011(*) 0.013(*)
-weighted $\phi_{L_{1}}(\bm{a})$ 0.002(**) 0.008(**) 0.002(**)
-weighted $\phi_{L_{2}}(\bm{a})$ 0.013(*) 0.015(*) 0.005(**)
+$\phi_{L_{1}}(\bm{a})$ 0.0001(***) 0.043(*) 0.011(*)
+$\phi_{L_{2}}(\bm{a})$ 0.0008(***) 0.026(*) 0.020(*)
+weighted $\phi_{L_{1}}(\bm{a})$ 0.0008(**) 0.004(**) 0.002(**)
+weighted $\phi_{L_{2}}(\bm{a})$ 0.002(*) 0.015(*) 0.006(**)
 (*) = significant at the $5\%$ significance level. (**) = significant at the $1\%$ significance level. (***) = significant at the $0.1\%$ significance level.
-It can be seen from Table that 5 $H_{0}$ is rejected in all cases.
+It can be seen from Table 5 that $H_{0}$ is rejected in all cases.
 In summary, there are significant differences between cleft and control subjects at all frames, with control subjects less asymmetric than cleft subjects overall.
-Since the test results indicate significant differences between the two groups, we would like to visualize the separation between the distributions of composite asymmetry scores of the two groups.
-Hence, dot plots similar to the stem leaf plot are created and given in Figures 4, 5, 6, 7 and 8.
-In these dot plots, composite asymmetry scores for both groups are plotted together at each frame.
-Figure 4 shows the dot plots for $\phi^{*}_{L_{1}}(\bm{a})$ defined in equation (3.9), whereas Figure 5 and 6 display the dot plots of $\phi_{L_{1}}(\bm{a})$ and $\phi_{L_{2}}(\bm{a})$ (equations (3.6) and (3.7)) respectively.
-According to these dot plots, substantial overlaps between cleft and control subjects can be seen.
-Moreover, the three composite asymmetry scores used in this section all have similar performances, as the separation between the two groups at each frame revealed by the dot plots are all similar.
-i First frame
-ii Middle frame
-iii Last frame
-Figure 4: Dot plots of $\phi^{*}_{L_{1}}(\bm{a})$ defined in equation (3.9) at the three frames. The top row of the graph is for cleft lip subjects whereas the bottom row is for controls.
-i First frame
-ii Middle frame
-iii Last frame
-Figure 5: Dot plots of $\phi_{L_{1}}(\bm{a})$ (3.6) at the three frames. The top row of the graph is for cleft lip subjects whereas the bottom row is for control subjects.
-i First frame
-ii Middle frame
-iii Last frame
-Figure 6: Dot plots of $\phi_{L_{2}}(\bm{a})$ (3.7) at the three frames. The top row of the graph is for cleft lip subjects whereas the bottom row is for control subjects.
-Figure 7 and 8 show the dot plots for weighted $\phi_{L_{1}}(\bm{a})$ and weighted $\phi_{L_{2}}(\bm{a})$ respectively.
-According to these figures, these two composite scores push outliers farther away, i.e., they are sensitive to outliers, especially for weighted $\phi_{L_{2}}(\bm{a})$.
-So the weighted $\phi_{L_{1}}(\bm{a})$ and weighted $\phi_{L_{2}}(\bm{a})$ emphasize which are more extreme in cleft lip group.
-i First frame
-ii Middle frame
-iii Last frame
-Figure 7: Dot plots of weighted $\phi_{L_{1}}(\bm{a})$ at the three frames. The top row of the graph is for cleft lip subjects whereas the bottom row is for control subjects.
-i First frame
-ii Middle frame
-iii Last frame
-Figure 8: Dot plots of weighted $\phi_{L_{2}}(\bm{a})$ at the three frames. The top row of the graph is for cleft lip subjects whereas the bottom row is for control subjects.
-Compare-then-combine approach using (3.4).
-The method described in Section 3.2 is used to select the landmarks which possess significant asymmetry information.
+5.2.2 Application of the Compare-then-combine Approach
+Feature selection The method described in Section 3.2 is used to select the landmarks which possess significant asymmetry information.
 $v_{j}$ is selected to be the two-sample t-statistics (equation (3.15)).
@@
 The total number of bootstrap iterations used was 10000.
-Boyett and Shuster (1977) suggest 17000 iterations to make sure the significance level is $1\%$.
-However, the results of using 10000 and 17000 iterations turned out to be the same for our smile data.
-The central pair, landmarks 6 and 8, correspond to $t$-values which exceed these critical points at all three frames.
-Thus, these two landmarks have the most significant information of asymmetry.
-Further, the simultaneous confidence interval related to landmarks 6 and 8 excludes 0, so that $H_{0}$ is rejected when use these landmarks.
-Further, most of the extreme $p$-values are obtained from Mann-Whitney U tests related to landmarks 6, 8 and two solo landmarks 7 and 19.
+Boyett and Shuster (1977) has suggested 17000 iterations to make sure the significance level is $1\%$.
+However, the results of using 10000 and 17000 iterations turned out to be the same for our Smile Data 1.
+The central pair, landmarks 6 and 8, corresponds to $t$-values which exceed these critical points at all three frames.
+Thus, these two landmarks carry the most significant information of asymmetry.
+Further, most of the extreme $p$-values obtained from Mann-Whitney U tests are related to landmarks 6, 8 and two solo landmarks 7 and 19.
 These tests are performed using $\phi_{L_{1}}$ and $\phi_{L2}$ computed on various subsets of landmarks.
@@
 Figure 4: Photo of a cleft patient with central pair landmarks and solo landmarks. This figure is reproduced from photo taken from the website https://www.nhs.uk/conditions/cleft-lip-and-palate/.
-We also applied the Welch version of the $t$-test (Welch, 1947), which accommodates different variances for the two groups.
+Meta-analysis Fisher’s method and Pearson’s method defined in equation (4.1) and (4.2) respectively are performed.
+Mann-Whitney U test is used to obtain individual left-sided $p$-values, $\tilde{p}_{j}$, for $j=1,\dots,35$, using $\bm{a}_{n}$, $n=1,\dots,25$.
+For Fisher’s method, we compare $P_{L}$ with the tail of $\chi_{70}^{2}$ distribution (as $J=35$) to obtain overall $p$-values.
+For the Pearson’s method, randomization tests with 10000 iterations are used to estimate the overall $p$-values.
+The values of $P_{L}$, $P_{R}$, $Q$ and $p$-values derived at the three frames are shown in Table 6.
+$H_{0}$ is rejected at all frames when we use Fisher’s method.
+For Pearson’s method, $H_{0}$ is rejected at the first and last frames.
+Comparing with Table 4 and 5, the $p$-values for the Fisher’s method are more extreme.
+This indicates that Fisher’s method is more sensitive to the departures of cleft group from control group.
+Table 6: The results of $P_{R}$ and $p$-values obtained from Fisher’s method and Pearson’s method at all three frames using the Smile Data 1.
+$P_{L}$ $P_{R}$ $Q$ Fisher’s method Pearson’s method
+First frame 206.65 12.26 206.65 0(***) $4\times 10^{-4}$(***)
+Middle frame 127.19 33.90 127.19 $3.51\times 10^{-5}$(***) 0.067
+Last frame 134.21 26.89 134.21 $6.18\times 10^{-6}$(***) 0.027(*)
+(*) = significant at the $5\%$ significance level. (**) = significant at the $1\%$ significance level. (***) = significant at the $0.1\%$ significance level.
+5.2.3 Welch’s t-test and Hotelling’s $T^{2}$ Test
+There are a few other tests we have applied and could be relevant for some future applications
+  • (a)
+    The Welch version of the $t$-test (Welch (1947), Lehmann and Romano (2005)), which accommodates different variances for the two groups.
 In this example, the results of the Welch $t$-test is found to be similar to the standard $t$-test (Table 4).
-A possible reason is that our sample size is small, so it results into similar outcome.
-Further, the multivariate Hotelling’s $T^{2}$ test on unsigned elementary feature vector $\bm{a}_{n}$, $n=1,\ldots,N$, is an alternative approach to the UIT.
-In this example, the test results fail to improve the outcomes of the univariate tests.
-It could be since a Hotelling’s $T^{2}$ test treats the features separately so it does not gain any power if the features point in the same direction; one-sided Hotelling’s $T^{2}$ could be appropriate approach, which we will pursue in future.
-In this example, the sample size is less than the dimension of $\bm{a}_{n}$, i.e.
+A possible reason is that our sample size is small, so it results in similar outcomes.
+  • (b)
+    Hotelling’s $T^{2}$ test (Mardia et al., 2024).
+We apply here Hotelling’s $T^{2}$ test on unsigned elementary feature vector $\bm{a}_{n}$, $n=1,\ldots,N$.
+In this example, we reject $H_{0}$ for the first two frames at $5\%$ significance level.
+The reason could be that a Hotelling’s $T^{2}$ test treats the features separately so it does not gain any power if the features point in the same direction; one-sided Hotelling’s $T^{2}$ could be an appropriate approach, which we will pursue in future.
+In this example of Smile Data 1, the sample size is less than the dimension of $\bm{a}_{n}$, i.e.
 $N<J$, so the covariance matrix is singular, which requires some modifications and will be dealt in future.
-We end this section by summarizing the main conclusions for this smile data:
-  • There are statistically significant differences between the two groups at all frames. So overall, control subjects are less asymmetric than the cleft lip subjects.
-  • Substantial overlaps can be found on dot plots in Figure 4, 5 and 6. This matches the expert’s expectation that some cleft lip subjects should be fairly close to normal subjects after surgeries.
+5.3 Visual Study Related to Three Composite Statistics
+(a) First frame
+(b) Middle frame
+(c) Last frame
+Figure 5: Plots for three composite asymmetry scores.
+From top to bottom, we have $\phi^{*}_{L_{1}}(\bm{a})$ (3.9) (in solid dots), $\phi_{L_{1}}(\bm{a})$ (3.6) (in triangles) and $\phi_{L_{2}}(\bm{a})$ (3.7) (in crosses) at the three frames computed on the Smile Data 1.
+Plot (a) is at first frame, while (b) is at middle frame, whereas (c) is at last frame.
+In plot (a), $\phi_{L_{1}}(\bm{a})$ and $\phi_{L_{2}}(\bm{a})$ are divided by 2.
+In plot (b) and (c), $\phi_{L_{2}}(\bm{a})$ is divided by 3.
+The first, third and fifth rows from top to bottom are cleft while others are control.
+Since the test results indicate significant differences between the two groups, we would like to visualize the distributions of $\phi^{*}_{L_{1}}(\bm{a})$, $\phi_{L_{1}}(\bm{a})$ and $\phi_{L_{2}}(\bm{a})$ defined in (3.9), (3.6) and (3.7) respectively between the two groups.
+Hence, dot plots similar to the stem leaf plot are created and given in Figure 5.
+In these dot plots, composite asymmetry scores for both groups are plotted together at each frame.
+According to these dot plots, substantial overlaps between cleft and control subjects can be seen.
+Moreover, the three scores used in this section all have similar performances, as the separation between the two groups at each frame revealed by the dot plots are all similar.
+Figure 6 shows the dot plots for weighted $\phi_{L_{1}}(\bm{a})$ and weighted $\phi_{L_{2}}(\bm{a})$ simultaneously.
+According to these figures, these two composite scores push outliers farther away from the rest of the samples, more so for weighted $\phi_{L_{1}}$.
+(a) First frame
+(b) Middle frame
+(c) Last frame
+Figure 6: Plots of two weighted asymmetry scores.
+From top to bottom, we have weighted $\phi_{L_{1}}(\bm{a})$ (in solid dots) and $\phi_{L_{2}}(\bm{a})$ (in triangles) at the three frames computed using the Smile Data 1.
+The first and third rows from top to bottom are cleft while others are control.
+Plot (a) is at first frame, while (b) is at middle frame, whereas (c) is at last frame.
+5.4 Summary of the Analyses
+We end this section by summarizing the main conclusions for these cleft lip data:
+  • (a)
+    There are statistically significant differences between the two groups at all frames. The control subjects are overall more symmetric than the cleft lip subjects.
+  • (b)
+    Substantial overlaps can be found on dot plots in Figure 5. This matches the expert’s expectation that some cleft lip subjects should be fairly close to normal subjects after surgeries.
+  • (c)
     The landmarks 6, 7, 8 and 19 (the ‘Y’-shape) are the most important in assessing asymmetries.
+6 Analyses of the Orthognathic Surgery Smile Data
+6.1 Details
+The orthognathic surgery smile data contains $N=22$ subjects who have undergone orthognathic surgery and their pre- and post-surgery data have both been collected.
+Figure 7 shows the face before (in the left) and after (in the right) the orthognathic surgery, where the view is from the right side of the face.
+It can be seen in the left of Figure 7, the jaw is somewhat protruded in comparison to the figure on right.
+Figure 7: View of the face from right before (left) and after (right) the orthognathic surgery.
+The axes used for the Smile Data 2 are exactly the same as the Smile Data 1 and the details are given in Section 2.1.
+The 16 landmarks on the nasal region are shown in left of Figure 8, while the right of Figure 8 shows the 10 landmarks on the lip.
+The landmark indices are summarized in Table 7.
+The scale of the Smile Data 2 is in mm.
+The same frames as the Smile Data 1 are used.
+Table 7: Indices of landmark pairs and solos in Figure 8 on the lip periphery and nasolabial region for the Smile Data 2.
+Landmark notation Indices in Figure 8
+pair $(k_{L},k_{R})$ (1,7), (2,6), (3, 5), (8,10), (11,12), (15,16)
+$k_{S}$ 4, 9, 13, 14
+There are two sets of landmarks we will focus on:
+  1. (a)
+    The 7 landmarks on the upper lip (landmarks 1 to 7 in Figure 8) and the 6 landmarks on the nasal region (landmarks 11 to 16 in the left of Figure 8).
+There are $K_{P}=5$ paired landmarks and $K_{S}=3$ solo landmarks, hence, $J=18$.
+  2. (b)
+    The 10 landmarks on the lip (landmarks 1 to 10 in Figure 8). There are $K_{P}=4$ paired landmarks and $K_{S}=2$ solo landmarks, hence, $J=14$.
+Figure 8: The left figure shows the 16 landmarks on the nasolabial region for a typical subject at the first frame using the pre-surgery data.
+The right figure shows the 10 landmarks on the lip periphery.
+The two figures are based on the Smile Data 2.
+These two sets of landmarks correspond to two aims:
+  1. (a)
+    Assess the affects of surgery, by comparing pre- and post-surgery data via paired tests (item (a) above).
+  2. (b)
+    Compare the subjects from Smile Data 2 to the control subjects from the Smile Data 1 where we use the same landmarks on the lip for both data (item (b) above).
+Some previous studies have been carried out, such as in Al-Hiyali et al.
+(2015), Xue et al.
+(2020), Xue et al.
+(2023) and Quast et al.
+(2023).
+These studies have focused on different facial expressions and the authors have taken average over measurements resulted from these expressions.
+However, by doing this, they would lose a lot of information.
+Our work is different from these previous studies as we attempt to construct asymmetry measurements over the lip and nasal region at different frames during the smile.
+The rest of this section will be divided into two parts corresponding to item (a) (effect of surgery, see Section 6.2) and item (b) (deviation from normal subjects, see Section 6.3) above.
+6.2 Test Results for Smile Data 2: Effect of Surgery
+The pre-surgery data is compared to the post-surgery data by using one-sided paired tests on the asymmetry measurements.
+Let $g=1,2$ denote the pre-surgery data and post-surgery data respectively.
+The absolute elementary feature vectors $\bm{a}_{n}^{(1)}$ and $\bm{a}_{n}^{(2)}$ are computed using the pre- and post-surgery data separately, for $n=1,\dots,22$.
+6.2.1 Application of the Combine-then-compare Approach
+The 13 landmarks on the upper lip and nasal region are used.
+Paired t-tests have been performed on composite asymmetry scores defined in equation (3.9) whereas Wilcoxon signed-rank tests have been performed on scores defined in (3.6), (3.7) respectively.
+Since the expert suspects that the asymmetries are increased after the surgery, so the one-sided alternative hypothesis becomes the distribution relates to pre-surgery data is stochastically smaller than that of post-surgery data.
+The Group 1 and Group 2 in (3.1) are changed to pre- and post-surgery data respectively.
+The test results on $\phi^{*}_{L_{1}}$ given in equation (3.9) are shown in Table 8.
+$H_{0}$ is rejected in middle and last frames.
+In other words, the asymmetries are increased after surgery when inspect at these two frames.
+Further, the $p$-value is the most extreme at the last frame, which indicates that the two groups are the most different at the last frame.
+Table 8: Mean, variance, t-values and $p$-values using $\phi^{*}_{L_{1}}(\bm{a})$ based on the Smile Data 2 (equation (3.9)).
+Pre Post One-sided paired test
+$\phi^{*}_{L_{1}}(\bm{a})$ mean sd mean sd $t$-values $p$-values
+First 15.69 5.64 18.47 8.49 -1.33 0.098
+Middle 17.53 5.92 22.25 10.89 -2.36 0.014(*)
+Last 17.72 5.45 22.77 9.52 -3.29 0.002(**)
+(*) = significant at the $5\%$ significance level. (**) = significant at the $1\%$ significance level. (***) = significant at the $0.1\%$ significance level.
+Table 9 displays the results of Wilcoxon signed-rank tests on $\phi_{L_{1}}(\bm{a})$ given by (3.6) and $\phi_{L_{2}}(\bm{a})$ given by (3.7).
+We reach the same outcomes as the paired t-tests: $H_{0}$ is rejected at the middle and last frames and the two groups are the most different at the last frame.
+Table 9: $p$-values from one-sided Wilcoxon signed-rank tests using $\phi_{L_{1}}(\bm{a})$ (3.6) and $\phi_{L_{2}}(\bm{a})$ (3.7) based on the Smile Data 2.
+Composite asymmetry score First frame Middle frame Last frame
+$\phi_{L_{1}}(\bm{a})$ 0.13 0.016(*) 0.002(**)
+$\phi_{L_{2}}(\bm{a})$ 0.088 0.007(**) 0.001(**)
+(*) = significant at the $5\%$ significance level. (**) = significant at the $1\%$ significance level. (***) = significant at the $0.1\%$ significance level.
+6.2.2 Application of the Compare-then-combine Approach
+Feature selection The procedures given in Section 3.2 are performed using pre- and post-surgery data, in order to select which landmarks contribute the most to the increment in asymmetries after the surgery.
+10000 bootstrap iterations are performed.
+The $25\%$, $20\%$, $15\%$ and $10\%$ critical values of $V$ are determined at the three frames separately.
+The corner pair, landmarks (1, 7) are important at the middle frame with respect to $25\%$, $20\%$ and $15\%$ critical values of $V$.
+The significance levels used here are larger than we used for Smile Data 1 in Section 5, which implies the features selected here are less important than the features in Smile Data 1.
+Meta-analysis Fisher’s method and Pearson’s method given in (4.1) and (4.2) are performed with Wilcoxon signed-rank tests as individual tests using $\bm{a}_{n}^{(1)}$ and $\bm{a}_{n}^{(2)}$, for $n=1,\dots,22$.
+$P_{L}$ defined in equation (4.1) is computed and compare with the tail of $\chi^{2}_{36}$ distribution (as $J=18$).
+As for $Q$, permutation tests are used to estimate the overall $p$-values.
+Table 10 shows the values of $P_{L}$, $P_{R}$, $Q$ and $p$-values for Fisher’s and Pearson’s method at the three frames.
+Table 10: The results of $P_{R}$ and $p$-values obtained from Fisher’s method and Pearson’s method at all three frames. Pre- and post-surgery data from the Smile Data 2 is used here.
+$P_{L}$ $P_{R}$ $Q$ Fisher’s method Pearson’s method
+First frame 45.52 20.04 45.52 0.13 0.45
+Middle frame 63.39 10.09 63.39 0.003(**) 0.085
+Last frame 64.20 11.53 64.20 0.003(**) 0.041(*)
+(*) = significant at the $5\%$ significance level. (**) = significant at the $1\%$ significance level. (***) = significant at the $0.1\%$ significance level.
+For the Fisher’s method, we obtain the same results as in the combine-then-compare approach: $H_{0}$ is rejected at middle and last frames.
+However, the $p$-values obtained in Fisher’s method are more extreme at the middle frame than those from combine-then-compare approach.
+For the Pearson’s method, we fail to reject $H_{0}$ at the first and middle frames.
+This indicates that the Pearson’s method may not be sensitive to changes in asymmetries before and after the surgery.
+6.3 Test Results for Smile Data 1 and Smile Data 2: Deviation from Control Subjects
+The Smile Data 2 is compared with the control subjects from the Smile Data 1.
+In Section 5, we only use 12 subjects which form a subset of the control group.
+The expert believes that more data should be used, thus, all 109 control subjects are used in this section.
+To make the comparisons available, a subset of 10 landmarks on the lip (Figure 8) are chosen from the 24 landmarks (see Figure 3), so that the Smile Data 2 and the Smile Data 1 have the same landmarks at the lip.
+Hence, the absolute elementary feature vectors $\bm{a}_{n}$ for $n=1,\dots,12$ are recomputed on the subjects from Smile Data 1 using the new landmarks.
+Pre- and post-surgery data are used separately.
+Hence, we have three groups in total, but we stick with two-sample tests.
+We scale the data so that both datasets have the same scale in cm.
+6.3.1 Application of the Combine-then-compare Approach
+The one-sided two-sample t-tests are performed on $\phi^{*}_{L_{1}}$ defined in (3.9) and the outcomes are shown in Table 11.
+Table 12 shows the means and variances of $\phi^{*}_{L_{1}}$ for the three groups of data.
+$H_{0}$ is rejected in almost all cases, except for the pre-surgery data at middle frame.
+This implies that in most of the cases, control subjects from the Smile Data 1 are more symmetric than subjects from the Smile Data 2 .
+Table 11: $p$-values from one-sided two-sample Welch’s $t$-test using $\phi_{L_{1}}^{*}$ defined in (3.9) to compare between Smile Data 2 subjects with normal subjects from the Smile Data 1.
+$\phi_{L_{1}}^{*}$ First frame Middle frame Last frame
+Pre-surgery 0.006(**) 0.073 0.031(*)
+Post-surgery 0.025(*) 0.013(*) 0.010(*)
+(*) = significant at the $5\%$ significance level. (**) = significant at the $1\%$ significance level. (***) = significant at the $0.1\%$ significance level.
+Table 12: Mean, standard deviation, t-values of $\phi^{*}_{L_{1}}(\bm{a})$ (equation (3.9)) for pre-, post-surgery subjects from Smile Data 2 and 12 control subjects from Smile Data 1.
+Pre Post Control
+$\phi^{*}_{L_{1}}(\bm{a})$ mean sd mean sd mean sd
+First 0.95 0.35 0.87 0.29 0.74 0.22
+Middle 1.17 0.41 1.30 0.52 1.03 0.35
+Last 1.22 0.34 1.32 0.46 1.07 0.33
+One-sided two-sample Mann-Whitney U tests have been performed using $\phi_{L_{1}}$ and $\phi_{L_{2}}$ defined in (3.6) and (3.7) respectively.
+The test results are shown in Table 13.
+$H_{0}$ is rejected in almost all cases except for pre-surgery data at the middle frame and post-surgery data at the first frame.
+Table 13: $p$-values obtained from one-sided two-sample Mann-Whitney U tests using $\phi_{L_{1}}$ in (equation (3.6)) and $\phi_{L_{2}}$ in (3.7) to compare between Smile Data 2 subjects with control subjects from the Smile Data 1.
+Pre Post
+$\phi_{L_{1}}$ $\phi_{L_{2}}$ $\phi_{L_{1}}$ $\phi_{L_{2}}$
+First 0.0044(**) 0.0050(**) 0.059 0.070
+Middle 0.062 0.060 0.0081(**) 0.014(*)
+Last 0.018(*) 0.040(*) 0.0030(**) 0.0037(**)
+(*) = significant at the $5\%$ significance level. (**) = significant at the $1\%$ significance level. (***) = significant at the $0.1\%$ significance level.
+However, the test results violate the expert’s observations: both pre- and post-surgery data are asymmetric at the first frame and the pre-surgery data is more asymmetric at this frame.
+We then compute the $\phi_{L_{1}}$ and $\phi_{L_{2}}$ with different weights for the solos, where the weights for the solos are the doubled unit lengths ($w_{j}=2$ in (3.6) and (3.7)) while we still keep the weights for landmark pairs to be unit length ($w_{j}=1$ in (3.6) and (3.7)).
+The test results from Mann-Whitney U tests are displayed in Table 14.
+$H_{0}$ is rejected for both groups at the first frame.
+Further, the $p$-values for pre-surgery group is more extreme, which indicates that the pre-surgery data are more asymmetric.
+This findings match the expert’s opinion.
+Note that changing the weights in this way do not affect the results for compare-then-combine approach.
+Table 14: $p$-values obtained from one-sided two-sample Mann-Whitney U tests using $\phi_{L_{1}}$ in (equation (3.6)) and $\phi_{L_{2}}$ in (3.7), where $w_{j}=1$ for pairs and $w_{j}=2$ for solos, to compare between Smile Data 2 subjects with control subjects from the Smile Data 1.
+Pre Post
+$\phi_{L_{1}}$ $\phi_{L_{2}}$ $\phi_{L_{1}}$ $\phi_{L_{2}}$
+First 0.0047(**) 0.0035(**) 0.026(*) 0.018(*)
+Middle 0.073 0.075 0.0055(**) 0.0073(**)
+Last 0.015(*) 0.042(*) 0.0018(**) 0.0026(**)
+(*) = significant at the $5\%$ significance level. (**) = significant at the $1\%$ significance level. (***) = significant at the $0.1\%$ significance level.
+The patterns of the $p$-values shown in Table 11, 13 and 14 are the same: the pre-surgery data has smaller $p$-values at the first frame, whereas the post-surgery data has more extreme $p$-values at the middle and last frames.
+6.3.2 Application of the Compare-then-combine Approach
+Feature selection The method described in Section 3.2 is used to select the landmarks which possess significant asymmetry information for the pre- and post-surgery data when compare with the normal subjects.
+$v_{j}$ is selected to be the two-sample t-statistics (equation (3.15)).
+$V$ defined in (3.16) is computed and its critical value at $5\%$ significance level is determined via bootstrap.
+The total number of bootstrap iterations used is 10000.
+The central pair, landmarks (3, 5), possesses t-values which exceed the critical values of $V$ in most of the cases for both pre- and post-surgery data.
+The solo landmark 9 also corresponds to t-values exceed the critical values of $V$ in some cases.
+Most of the extreme p-values obtained from Mann-Whitney U tests pertain to the central pair 3, 5 and two solos, 4 and 9, where $\phi_{L_{1}}$ and $\phi_{L_{2}}$ are used.
+Landmark pair (3, 5) is the same as landmark pair (6, 8) and solo landmarks 4 and 9 are the same as landmarks 7 and 19 in Figure 4, so we recover the ‘Y’-shape as in Section 5.
+Meta-analysis The Fisher’s and Pearson’s methods given in (4.1) and (4.2) are executed to compare the subjects from Smile Data 2 with control subjects from the Smile Data 1.
+$P_{R}$ defined in equation (4.1) is computed at the three frames separately and compared with the tail of $\chi^{2}_{28}$ (as $J=14$).
+Pre-surgery data Table 15 shows the results at the three frames which are obtained by comparing pre-surgery data with control subjects.
+Comparing with Table 11 and 14, $H_{0}$ is also rejected at the middle frame for the Fisher’s method and the $p$-values are more extreme at the first and last frames.
+The reason could be that the means of some asymmetry features are different for the two groups.
+However, these features have small magnitudes, so they do not contribute significantly towards the values of composite scores.
+Consequently, they do not affect the test results on the composite scores, but impact the outcomes of Fisher’s method.
+Nevertheless, for the Pearson’s method, the results are the same as in Table 11 and 14 and the magnitudes of $p$-values are similar as well.
+Table 15: The results of $P_{R}$ and $p$-values obtained from Fisher’s method and Pearson’s method at all three frames.
+Pre-surgery data from the Smile Data 2 is compared with the control subjects from the Smile Data 1 here.
+$P_{L}$ $P_{R}$ $Q$ Fisher’s method Pearson’s method
+First frame 9.54 67.13 67.13 $4.63\times 10^{-5}$(***) 0.003(**)
+Middle frame 22.25 46.44 46.44 0.016(*) 0.11
+Last frame 17.16 51.90 51.90 0.0039(**) 0.038(*)
+(*) = significant at the $5\%$ significance level. (**) = significant at the $1\%$ significance level. (***) = significant at the $0.1\%$ significance level.
+Post-surgery data Table 16 shows the results obtained by comparing post-surgery data and control subjects.
+We obtain the same test results as in Table 11 and 14.
+The magnitudes of $p$-values for Pearson’s method are similar with the $p$-values obtained for $\phi_{L_{1}}$ and $\phi_{L_{2}}$ at the middle and last frames, whereas the $p$-value is decreased at the first frame for Pearson’s method.
+For Fisher’s method, the $p$-values are more extreme.
+Table 16: The results of $P_{R}$ and $p$-values obtained from Fisher’s method and Pearson’s method at all three frames.
+Post-surgery data from the Smile Data 2 is compared with the control subjects from the Smile Data 1 here.
+$P_{L}$ $P_{R}$ $Q$ Fisher’s method Pearson’s method
+First frame 14.26 67.61 67.61 $3.98\times 10^{-5}$(***) 0.003(**)
+Middle frame 10.83 70.64 70.64 $1.51\times 10^{-5}$(***) 0.004(**)
+Last frame 13.64 63.86 63.86 0.00013(***) 0.009(**)
+(*) = significant at the $5\%$ significance level. (**) = significant at the $1\%$ significance level. (***) = significant at the $0.1\%$ significance level.
+6.4 Summary of the Analyses
+We end the section by summarizing our findings on the orthognathic surgery data:
+  • (a)
+    The asymmetries are increased after the surgery at middle and last frames.
+  • (b)
+    The subjects contained in the Smile Data 2 are more asymmetric than the control subjects from the Smile Data 1, except for the pre-surgery data at the middle frame.
+  • (c)
+    The central pair and solo landmarks, i.e. the ‘Y’-shape in Figure 4, carry important asymmetry information when we compare the subjects in Smile Data 2 with control subjects from Smile Data 1.
 7 Discussion
-We have presented our work for pre-registered data, but it can be extended when the landmark data is not pre-registered as pointed out in Section 4.
-Further, these revised measures of asymmetry can be applied and analyzed in a straightforward way.
-Further, we have used only size-and-shape framework but it can be extended easily to include for example, similarity shape.
-We have used absolute values ($a_{j}$) on coordinate elementary features as in equation (2.5) which is of main interest if one is looking for the extent of asymmetry as in our case for the smile data.
-However, there maybe some other applications where raw $d_{j}$ may be important than $a_{j}$; our work easily applies to that situation.
-In this paper we have concentrated on smile example with three fixed frames.
-Our future statistical work will be extended to dynamic shape analysis which will allow to study trajectories of the smile.
-Some initial work for a single group of a smile data has been introduced in Mardia et al.
-(2018) and Bookstein et al.
-(2024).
-We have concentrated on the cleft lip subjects but there are other medical applications such as in facial reconstruction related to trauma surgery where the lips form significant area for surgery.
-We have used univariate methods but we can use Hotelling’s $T^{2}$ in place of t-statistics, for example, though it requires some modifications.
-An one-sided alternative seems natural according to equation (3.1) and in future we would consider one-sided multivariate tests.
-Nevertheless, it requires more care to set up than in the one-dimensional case.
-Further, the scenario of sample size is less than the dimension of the data vector has also to be taken into account.
-Acknowledgements
-The authors would like to thank the School of Mathematics, University of Leeds for supporting the smile project. Kanti Mardia acknowledges the Leverhulme Trust for the Emeritus Fellowship.
-Appendix A Estimation of midplane using Procrustes analysis
-Let $\bm{n}\in\mathbb{R}^{M}$ denote the unit normal vector of an arbitrary plane $\mathcal{P}^{\prime}$.
-Let $X_{n}\in\mathbb{R}^{K\times M}$, $n=1,\dots,N$ denote the observed configurations.
-The reflection of $X_{n}$ about $\mathcal{P}^{\prime}$, $X^{\text{(refl)}}_{n}$, is given as
-$X^{\text{(refl)}}_{n}=X_{n}H,$
-where $H=I_{M}-2\bm{n}\bm{n}^{T}$ is the Householder matrix and $I_{M}$ is the $M\times M$ identity matrix. Then the estimation process used in basis registration is given as the following:
-  1. 1.
-    Form an augmented dataset: $\{X_{n},X^{\text{(refl)}}_{n}\}_{n=1}^{N}$.
-  2. 2.
-    Apply GPA on this augmented dataset. Denote the fitted configurations as $X_{n}^{(\text{GPA})}$ for $n=1,\dots,N$.
-  3. 3.
-    Compute the Procrustes mean shape. This mean shape will be bilaterally symmetric with midplane $\mathcal{P}^{\prime}$.
-  4. 4.
-    Rotate the mean shape outside and within $\mathcal{P}^{\prime}$ until the new midplane and the basis are meaningful.
-  5. 5.
-    Carry out ordinary Procrustes analysis between each $X_{n}^{(\text{GPA})}$ and the rotated mean shape.
-When axis registration is considered, we only rotate the sample mean shape outside midplane.
-Further, the last step, step 5, is ignored.
-Further details are given for this procedure in Mardia et al.
-(2000).
+We have presented our asymmetry measures for the registered Smile Data.
+But these asymmetry measures are available to any registered dataset.
+Further, we have used only a size-and-shape framework but the composite measures can be extended easily to include for example, similarity shape, since once we have registered the shapes, the asymmetry measures are applicable after allowing for scaling.
+We have used absolute values ($a_{j}$) on coordinate elementary features as in equation (2.5) which is of main interest if one is looking for the extent of asymmetry as in our cases for the illustrative smile examples.
+However, there maybe some other applications where raw $d_{j}$ may be important than $a_{j}$, such as when analyzing the trajectories of the landmarks; our work easily applies to that situation.
+In this paper we have concentrated on the illustrative smile examples with three fixed frames and our future statistical work will be extended to dynamic shape analysis which will allow us to study trajectories of the smile.
+In our case, the number of landmarks $K$ is small but when $K$ is large we can study the distribution of asymmetry measure $\phi_{L_{2}}$ under isotropic normality and various tests can be then performed.
+The work leads to understanding for the sum of weighted $\chi^{2}$-distributions.
+Mardia and Wu (2025) studies this approach.
+We have used mostly univariate methods except for the Hotelling’s $T^{2}$ in Section 5, though in our case, the sample size is less than the dimension of the elementary feature vector so it is not appropriate as Hotelling’s $T^{2}$ is singular.
+Even in the case when the sample size is large enough, the one-sided alternative seems natural according to equation (3.1).
+Such cases are also part of our future investigation.
+We note that the final aim of our work is to provide in future a smile index based on these asymmetry measures to be used as a standard index in clinics.
+Data Availability Statement
+The codes and data for this paper are available at the following website: https://github.com/XW-2025-hub/JRSS-C-Supplementary-Materials.
+Acknowledgments
+The authors would like to thank the School of Mathematics, and the School of Dentistry University of Leeds for supporting the smile project; we are also grateful to Fred Bookstein for collaborating with this project.
+The authors would also like to thank Sofya Titarenko for her helpful comments.
+Kanti Mardia acknowledges the Leverhulme Trust for the Emeritus Fellowship.